### v5

# N-dimensional Convex Obstacle Avoidance using Hybrid Feedback Control (Extended version)

Abstract
This paper addresses the autonomous robot navigation problem in a priori unknown $n$-dimensional environments containing convex obstacles of arbitrary shapes and sizes. We propose a hybrid feedback control scheme that guarantees safe and global asymptotic convergence of the robot to a predefined target location. The proposed control strategy relies on a switching mechanism allowing the robot to operate either in the move-to-target mode or the obstacle-avoidance mode, based on its proximity to the obstacles and the availability of a clear straight path between the robot and the target. In the obstacle-avoidance mode, the robot is constrained to move within a two-dimensional plane that intersects the obstacle being avoided and the target, preventing it from retracing its path. The effectiveness of the proposed hybrid feedback controller is demonstrated through simulations in two-dimensional and three-dimensional environments.
I Introduction
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One commonly used navigation technique is the Artificial Potential Field (APF) approach [1], where a combination of attractive and repulsive vector fields guides the robot safely to the target location. However, this approach faces challenges with certain obstacle arrangements, leading to undesired stable local minima. The Navigation Function (NF) approach [2, 3] is effective in sphere world environments, addressing the local minima issue by limiting the repulsive field’s influence around the obstacles by means of a properly tuned parameter. However, this method ensures almost [^1] Almost global convergence refers to the convergence from all initial conditions except a set of zero Lebesgue measure. global convergence of the robot to the target location. To apply the NF approach to environments with general convex and star-shaped obstacles, one has to use diffeomorphic mappings from [2] and [4], which require global knowledge of the environment for implementation. In [5], the authors extended the NF approach to handle environments with convex obstacles with smooth boundaries that meet certain curvature conditions. However, this approach is limited to obstacles, which are not too flat and not too close to the target. In [6], for the case of ellipsoidal worlds, the authors removed the flatness limitation in [5], by providing a controller design, which locally transforms the region near the obstacle into a spherical region by using the Hessian information. However, similar to [5], it is assumed that the entire shape of the obstacle becomes known when the robot visits its neighbourhood.
Other approaches, such as [7, 8, 9], rely on the control Lyapunov function (CLF) and the control barrier function (CBF) to design feedback control laws that achieve (simultaneously) convergence to a target set and avoidance of an unsafe set. In [8], the authors proposed a CBF-based method for multi-robot navigation in two-dimensional environments in the presence of circular obstacles. In [9], a comparative analysis between the APF-based and CBF-based approaches has been provided. However, the work in [10] and [11] demonstrates that the CLF-CBF-based navigation approach, similar to the NF approach, suffers from the undesired equilibria problem and provides (at best) almost global asymptotic stability guarantees in sphere worlds.
In [12], the authors proposed a feedback controller based on Nagumo’s theorem [13, Theorem 4.7] for autonomous navigation in environments with general convex obstacles. The forward invariance of the obstacle-free space is achieved by projecting the ideal velocity control vector, which points towards the target, onto the tangent cone at the obstacle boundary whenever it points towards the obstacle. This approach was extended in [14] to guarantee almost global asymptotic stability of the target location in a priori unknown environments containing strongly convex obstacles.
In [15], the authors proposed a purely reactive autonomous navigation approach based on separating hyperplanes for robots operating in environments cluttered with unknown but sufficiently separated convex obstacles with smooth boundaries which satisfy curvature conditions similar to [14]. This approach was extended in [16] to address partially known environments with non-convex obstacles, where the robot has geometric information about the non-convex obstacles but lacks precise knowledge of their locations in the workspace.
The approaches discussed above provide, at best, almost global convergence guarantees due to the undesired equilibria that are generated when using continuous time-invariant vector fields [17]. This can be resolved by introducing discontinuities in the control, as shown in [18, 19, 20, 21, 22, 23, 24].
In [18] and [19], hybrid control methods are employed to achieve robust global asymptotic stabilization in $\mathbb{R}^{2}$ for robots navigating towards a target location while avoiding collision with a single spherical obstacle. The approach in [18] has been extended in [20] to steer a group of planar robots in formation toward the source of an unknown but measurable signal while avoiding a single obstacle. In [21], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighbourhoods of unsafe isolated points. In [22], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
In [23], the authors proposed a discontinuous feedback-based autonomous navigation scheme for nonholonomic robots operating in two-dimensional environments with non-convex obstacles, subject to restrictions on inter-obstacle arrangements. In [24], a discontinuous feedback control law was introduced for autonomous robot navigation in partially known two-dimensional environments. When encountering a known obstacle, the control vector aligns with the negative gradient of the navigation function, whereas near unknown obstacles, the robot follows the boundary, using local curvature information.
In our earlier works [25] and [26], we proposed a hybrid feedback control strategy for autonomous robot navigation in two-dimensional environments with arbitrarily-shaped convex and non-convex obstacles. In this paper, we extend this approach by proposing a hybrid control strategy with global asymptotic stability guarantees for autonomous navigation in a priori unknown $n$-dimensional environments containing arbitrarily-shaped convex obstacles, including those with non-smooth boundaries. The main contributions of the proposed research work are as follows:
  1. 1.
    Global asymptotic stability: The proposed autonomous navigation solution guarantees global asymptotic stabilization of the target location in unknown environments with convex obstacles of arbitrary shapes. Note that the few existing results in the literature achieving such strong stability results are of a hybrid type and are restricted to environments with ellipsoidal obstacles [22].
  2. 2.
    $n$-dimensional convex obstacles: The proposed hybrid feedback controller is applicable to $n$-dimensional workspaces containing convex obstacles of arbitrary shapes and sizes. In contrast, the autonomous navigation schemes in [23] and [24] are limited to two-dimensional environments, while the methods in [6] and [12] apply only to $n$-dimensional environments with ellipsoidal obstacles. The navigation approaches in [14] and [15] are restricted to environments with strongly convex obstacles.
  3. 3.
    Arbitrary interobstacle arrangements: There are no restrictions on the arrangement of obstacles, unlike those imposed in [23, Assumption 10] and [22, Theorem 2], except for the widely accepted mild condition in Assumption 9, which states that the robot can pass between any two obstacles while maintaining a positive distance.
  4. 4.
    Applicable in a priori unknown environments: Unlike the approaches in [5] and [6], which require global information about the obstacles, the proposed autonomous navigation method relies solely on range scanners and does not require a priori global knowledge of the environment (sensor-based technique).
The remainder of the paper is organized as follows. Section II introduces the notations and preliminaries used throughout the paper. The problem formulation is presented in Section III, followed by the proposed hybrid control algorithm in Section IV. Stability and safety guarantees of the navigation control scheme are discussed in Section V. In Section VI, the hybrid feedback control law is modified to ensure a monotonic decrease in distance to the target in sphere worlds. Section VII outlines the implementation procedure for the obstacle avoidance algorithm in a priori unknown environments. Simulation results are provided in Section VIII to demonstrate the algorithm’s effectiveness, and concluding remarks are given in Section IX.
II Notations and Preliminaries
II-A Notations
The sets of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We identify vectors using bold lowercase letters. The Euclidean norm of a vector $\mathbf{p}\in\mathbb{R}^{n}$ is denoted by $\lVert\mathbf{p}\rVert$, and an Euclidean ball of radius $r\geq 0$ centered at $\mathbf{p}$ is represented by $\mathcal{B}_{r}(\mathbf{p})=\{\mathbf{q}\in\mathbb{R}^{n}|\lVert\mathbf{q}-\mathbf{p}\rVert\leq r\}.$ The set of $n-$dimensional unit vectors is given by $\mathbb{S}^{n-1}=\{\mathbf{p}\in\mathbb{R}^{n}|\lVert\mathbf{p}\rVert=1\}.$ The identity matrix of order $n$ is denoted by $\mathbf{I}_{n}$. The $n-$dimensional Special Orthogonal group is denoted by $SO(n):=\{\mathbf{R}\in\mathbb{R}^{n\times n}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{n},\det(\mathbf{R})=1\}.$
For two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$, the relative complement of $\mathcal{B}$ with respect to $\mathcal{A}$ is denoted by $\mathcal{A}\setminus\mathcal{B}=\{\mathbf{a}\in\mathcal{A}|\mathbf{a}\notin\mathcal{B}\}$. The symbols $\partial\mathcal{A},\mathcal{A}^{\circ}$, $\mathcal{A}^{c}$ and $\bar{\mathcal{A}}$ represent the boundary, interior, complement and the closure of the set $\mathcal{A}$, respectively, where $\partial\mathcal{A}=\bar{\mathcal{A}}\backslash\mathcal{A}^{\circ}$. The cardinality of a set $\mathcal{A}$ is denoted by $\mathbf{card}(\mathcal{A})$. The Minkowski sum of the sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by $\mathcal{A}\oplus\mathcal{B}=\{\mathbf{a}+\mathbf{b}|\mathbf{a}\in\mathcal{A},\mathbf{b}\in\mathcal{B}\}$. The dilated version of a set $\mathcal{A}\subset\mathbb{R}^{n}$ with $r\geq 0$ is represented by $\mathcal{D}_{r}(\mathcal{A})=\mathcal{A}\oplus\mathcal{B}_{r}(\mathbf{0})$. The $r-$neighbourhood of a set $\mathcal{A}$ is denoted by $\mathcal{N}_{r}(\mathcal{A})=\mathcal{D}_{r}(\mathcal{A})\setminus\mathcal{A}^{\circ}$, where $r$ is a strictly positive scalar.
II-B Projection on a set
Given a closed set $\mathcal{A}\subset\mathbb{R}^{n}$ and a point $\mathbf{x}\in\mathbb{R}^{n}$, the Euclidean distance of $\mathbf{x}$ from the set $\mathcal{A}$ is evaluated as
$d(\mathbf{x},\mathcal{A})=\underset{\mathbf{q}\in\mathcal{A}}{\min}\lVert\mathbf{x}-\mathbf{q}\rVert.$ (1)
If $\mathcal{A}$ is convex, the unique closest point to any $\mathbf{x}\in\mathbb{R}^{n}$ on $\mathcal{A}$ is denoted by $\Pi(\mathbf{x},\mathcal{A})$ and is defined as
$\Pi(\mathbf{x},\mathcal{O}_{i}):=\underset{\mathbf{q}\in\mathcal{A}}{\arg\min}\lVert\mathbf{x}-\mathbf{q}\rVert.$ (2)
II-C Geometric subsets of $\mathbb{R}^{n}$
II-C1 Line
The line passing through two points $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{q}\in\mathbb{R}^{n}\setminus\{\mathbf{p}\}$ is given by
$\mathcal{L}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{x}=\lambda\mathbf{p}+(1-\lambda)\mathbf{q},\lambda\in\mathbb{R}\}.$ (3)
II-C2 Line segment
The line segment joining two points $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{q}\in\mathbb{R}^{n}\setminus\{\mathbf{p}\}$ is given by
$\mathcal{L}_{s}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{x}=\lambda\mathbf{p}+(1-\lambda)\mathbf{q},\lambda\in[0,1]\}.$ (4)
II-C3 Plane
Given two linearly independent vectors $\mathbf{p},\mathbf{q}\in\mathbb{R}^{n}$, the plane containing all vectors that can be expressed as a linear combination of $\mathbf{p}$ and $\mathbf{q}$ is given by
$\mathcal{P}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\exists k_{1}\in\mathbb{R},\exists k_{2}\in\mathbb{R},\mathbf{x}=k_{1}\mathbf{p}+k_{2}\mathbf{q}\}$ (5)
II-C4 Hyperplane
The hyperplane passing through $\mathbf{p}\in\mathbb{R}^{n}$ and orthogonal to $\mathbf{q}\in\mathbb{R}^{n}\backslash\{\mathbf{0}\}$ is given by
$\mathcal{H}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{q}^{\intercal}(\mathbf{x}-\mathbf{p})=0\}.$ (6)
The hyperplane divides the Euclidean space $\mathbb{R}^{n}$ into two half-spaces i.e., a closed positive half-space $\mathcal{H}_{\geq}(\mathbf{p},\mathbf{q})$ and a closed negative half-space $\mathcal{H}_{\leq}(\mathbf{p},\mathbf{q})$ which are obtained by substituting ‘$=$’ with $\geq$’ and $\leq$’ respectively, in the right-hand side of (6). We also use the notations $\mathcal{H}_{>}(\mathbf{p},\mathbf{q})$ and $\mathcal{H}_{<}(\mathbf{p},\mathbf{q})$ to denote the open positive and the open negative half-spaces such that $\mathcal{H}_{>}(\mathbf{p},\mathbf{q})=\mathcal{H}_{\geq}(\mathbf{p},\mathbf{q})\backslash\mathcal{H}(\mathbf{p},\mathbf{q})$ and $\mathcal{H}_{<}(\mathbf{p},\mathbf{q})=\mathcal{H}_{\leq}(\mathbf{p},\mathbf{q})\backslash\mathcal{H}(\mathbf{p},\mathbf{q})$.
II-C5 Cylinder
Given $\mathbf{p}\in\mathbb{R}^{n}$ and $w>0$, the cylinder with width $w$ and the line segment $\mathcal{L}_{s}(\mathbf{p},\mathbf{0})$ as its axis is denoted by
$\mathcal{C}_{\mathcal{L}}(\mathbf{p},w):=\mathcal{D}_{w}(\mathcal{L}_{s}(\mathbf{p},\mathbf{0}))\cap\mathcal{H}_{\geq}(\mathbf{p},-\mathbf{p})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{p})$
II-D Parallel projection operator
Given two orthonormal vectors $\mathbf{p},\mathbf{q}\in\mathbb{S}^{n-1}$, the parallel projection operator is defined as
$\mathbf{P}(\mathbf{p},\mathbf{q})=\mathbf{p}\mathbf{p}^{\top}+\mathbf{q}\mathbf{q}^{\top}.$
For any vector $\mathbf{s}\in\mathbb{R}^{n}$, the vector $\mathbf{s}^{\prime}$, where $\mathbf{s}^{\prime}=\mathbf{P}(\mathbf{p},\mathbf{q})\mathbf{s}$, is the projection of $\mathbf{s}$ onto the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$. Additionally, it can be verified that $\mathbf{s}^{\top}\mathbf{s}^{\prime}\geq 0$, indicating that the angle between $\mathbf{s}$ and $\mathbf{s}^{\prime}$ is either acute or $90^{\circ}$.
II-E n-Dimensional rotation matrix
Consider two orthonormal vectors $\mathbf{p},\mathbf{q}\in\mathbb{S}^{n-1}$ and an angle $\theta\in[0,2\pi)$. The rotation matrix $R(\theta,\mathbf{p},\mathbf{q})$ is constructed as follows:
$R(\theta,\mathbf{p},\mathbf{q}):=\mathbf{I}_{n}+\sin(\theta)\mathbf{Sk}(\mathbf{p},\mathbf{q})+(1-\cos(\theta))(\mathbf{Sk}(\mathbf{p},\mathbf{q}))^{2},$
where the skew-symmetric matrix $\mathbf{Sk}(\mathbf{p},\mathbf{q})$ is defined by
$\mathbf{Sk}(\mathbf{p},\mathbf{q})=\mathbf{q}\mathbf{p}^{\top}-\mathbf{p}\mathbf{q}^{\top}.$
The operator $R(\theta,\mathbf{p},\mathbf{q})$ performs a rotation by the angle $\theta$ in the plane spanned by $\mathbf{p}$ and $\mathbf{q}$. Rotation angles are considered positive if they are performed from the vector $\mathbf{p}$ to the vector $\mathbf{q}$.
II-F Hybrid system framework
A hybrid dynamical system [27] is represented using differential and difference inclusions for the state $\mathbf{\xi}\in\mathbb{R}^{n}$ as follows:
$\begin{cases}\begin{matrix}\mathbf{\dot{\xi}}\in\mathbf{F}(\mathbf{\xi}),&\mathbf{\xi}\in\mathcal{F},\\ \mathbf{\xi}^{+}\in\mathbf{J}(\mathbf{\xi}),&\mathbf{\xi}\in\mathcal{J},\end{matrix}\end{cases}$ (7)
[…]
A subset $\mathbb{T}\subset\mathbb{R}_{\geq}\times\mathbb{N}$ is a hybrid time domain if it is a union of a finite or infinite sequence of intervals $[t_{j},t_{j+1}]\times\{j\},$ where the last interval (if existent) is possibly of the form $[t_{j},T)$ with $T$ finite or $T=+\infty$. The ordering of points on each hybrid time domain is such that $(t,j)\preceq(t^{\prime},j^{\prime})$ if $t\leq t^{\prime}$ and $j\leq j^{\prime}$. A hybrid solution $\phi$ is maximal if it cannot be extended, and complete if its domain dom $\phi$ (which is a hybrid time domain) is unbounded.
III Problem Formulation
Let $\mathcal{W}$ be a closed subset of the $n-$dimensional Euclidean space that bounds the workspace. The workspace $\mathcal{W}$ consists of finite number of compact, convex obstacles $\mathcal{O}_{i}$, $i\in\{1,\ldots,b\},b\in\mathbb{N}$. We define obstacle $\mathcal{O}_{0}:=(\mathcal{W}^{\circ})^{c}$ as the complement of the interior of the workspace. Collectively, the obstacle-occupied workspace is denoted by $\mathcal{O}_{\mathcal{W}}=\bigcup_{i\in\mathbb{I}}\mathcal{O}_{i}$, where $\mathbb{I}=\{0,\ldots,b\}.$
The robot is represented by a $n-$dimensional sphere with radius $r\geq 0$ and center $\mathbf{x}$. It is equipped with a range sensor that can identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ which contains the locations on the boundaries of nearby obstacles, provided there is a clear line of sight to the center of the robot, up to a certain sensing range $R_{s}$, as shown in Fig. 1. The set ð$\eth_{\mathcal{W}}$ is defined as follows:
ð$\eth_{\mathcal{W}}^{\mathbf{x}}:=\{\mathbf{p}\in\partial\mathcal{O}_{\mathcal{W}}|\|\mathbf{p}-\mathbf{x}\|\leq R_{s},\mathcal{L}_{s}(\mathbf{p},\mathbf{x})\cap\mathcal{O}_{\mathcal{W}}=\mathbf{p}\}.$ (8)
Figure 1: Illustration of the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (8).
To maintain the local convexity of the obstacle boundaries, we impose the following workspace feasibility assumption.
Assumption 1.
The minimum separation between any pair of obstacles should be greater than or equal to $2(r+\delta)$ i.e., for all $i,j\in\mathbb{I},i\neq j,$ one has
$d(\mathcal{O}_{i},\mathcal{O}_{j}):=\underset{\mathbf{p}\in\mathcal{O}_{i},\mathbf{q}\in\mathcal{O}_{j}}{\text{min}}\|{\mathbf{p}-\mathbf{q}}\|\geq 2(r+\delta),$ (9)
where $\delta>0$.
We then pick an arbitrarily small value $r_{s}\in(0,\delta)$ as the minimum distance that the robot should maintain to any obstacle. Therefore, the obstacle-free workspace with respect to the center of the robot is defined as follows:
$\mathcal{W}_{r_{a}}:=\mathcal{W}\setminus(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ},$ (10)
where $r_{a}=r+r_{s}$. According to Assumption 9, the set $\mathcal{W}_{r_{a}}$ is pathwise connected and $\mathbf{x}\in\mathcal{W}_{r_{a}}\iff\mathcal{B}_{r_{a}}(\mathbf{x})\subset\mathcal{W}\setminus\mathcal{O}_{\mathcal{W}}^{\circ}$.
The robot is governed by a single integrator dynamics
$\dot{\mathbf{x}}=\mathbf{u},$ (11)
where $\mathbf{u}\in\mathbb{R}^{n}$ is the control input. Given a workspace that satisfies Assumption 9, and given that the robot can identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ (8), the task is to design a feedback control law $\mathbf{u}$ to guarantee the following properties:
  1. 1.
    Safety: the obstacle-free workspace $\mathcal{W}_{r_{a}}$ with respect to the center of the robot is forward invariant,
  2. 2.
    Global asymptotic stability: any target location $\mathbf{x^{d}}\in(\mathcal{W}_{r_{a}})^{\circ}$ is a globally asymptotically stable equilibrium for the closed-loop system. Without loss of generality, we will consider $\mathbf{x^{d}}=\mathbf{0}$.
Figure 2: Robot trajectory moving along the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ while avoiding an obstacle.
IV Hybrid Control For Obstacle Avoidance
The proposed hybrid feedback controller operates in two different modes based on the mode indicator $m\in\mathbb{M}:=\{0,1\}$. In the move-to-target mode $(m=0)$, the control input vector directs the robot along a straight line toward the target location. When the robot is in the neighbourhood of an obstacle that obstructs its direct path (i.e., the distance from the robot’s centre to the closest point on the obstacle is less than or equal to $\gamma\in(0,\delta-r_{s})$), the control law switches to the obstacle-avoidance mode $(m=1)$. In contrast to the two-dimensional case, where the robot can only move clockwise or counterclockwise around an obstacle, the three-dimensional environment offers an infinite number of safe paths around the obstacles. Thus, designing an obstacle-avoidance strategy that maintains the continuity of the control input vector and prevents the robot from retracing its path is crucial. The proposed strategy achieves this by confining the robot’s motion to a plane that passes through both the target location and the interior of the obstacle being avoided. The robot is guided along the obstacle’s boundary until it reaches a point where the obstacle no longer blocks the direct line-of-sight to the target location, as shown in Fig. 2.
IV-A Hybrid control design
The proposed hybrid control $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ is given as
$\mathbf{u}(\xi)=-\kappa_{s}(1-m)\mathbf{x}+\kappa_{r}m\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a}),$ (12a)
$\underbrace{\begin{matrix}\mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}\;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}}\quad\quad\underbrace{\begin{bmatrix}\mathbf{h^{+}}\\ \mathbf{a^{+}}\\ m^{+}\\ {s^{+}}\end{bmatrix}\begin{matrix}\;\in\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s),\end{matrix}}_{(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{J}}$ (12b)
where $\kappa_{s}>0$, $\kappa_{r}>0$, and the composite state vector $\xi:=(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ The variable $\mathbf{h}$ denotes the hit point, which is the location where the robot switches from the move-to-target mode to the obstacle-avoidance mode. The vector $\mathbf{a}\in\mathbb{S}^{n-1}$ is instrumental for the construction of the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, used in (12a). The scalar variable $s\in\mathbb{R}_{\geq 0}$ allows the robot to switch once from the obstacle-avoidance mode to the move-to-target mode only when it is initialized in the obstacle-avoidance mode. Details of this switching process are provided later in Section IV-B. The sets $\mathcal{F}$ and $\mathcal{J}$ are the flow and jump sets related to different modes of operation, respectively, whose constructions are provided in Section IV-B. The update law $\mathbf{L}$, which allows the robot to update the values of the variables $\mathbf{h}$, $\mathbf{a}$, $m$ and $s$ based on the current location of the robot with respect to the obstacle being avoided and the target location, will be designed later in Section IV-C. Next, we provide the design of the vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathbb{R}^{n}$. The vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, used in (12a), is defined as
$\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})=\left[\eta(\mathbf{x})+(1-|\eta(\mathbf{x})|)\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\right]\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},$ (13)
where $\hat{\mathbf{h}}$ denotes unit vector in the direction of $\mathbf{h}$, and $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. The vector $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ represents the point on the obstacle-occupied workspace closest to $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 9, the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighbourhood of these obstacles. On the other hand, since obstacle $\mathcal{O}_{0}$, where $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$, is non-convex, there may be some locations in the $(r_{a}+\gamma)-$neighbourhood of the obstacle $\mathcal{O}_{0}$ for which the uniqueness of the closest point from the robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 2, the design of the flow sets and the jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
Note that the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ are updated when the robot switches from the move-to-target mode to the obstacle-avoidance mode using the update law $\mathbf{L}(\xi)$, whose design is provided later in Section IV-C. It is ensured that $\mathbf{a}^{\top}\mathbf{h}=0$, which allows us to define the operators $\mathbf{P}$ and $\mathbf{R}$, used in (13), as discussed next.
Since $\mathbf{h}$ and $\mathbf{a}$ are ensured to be orthogonal to each other, the parallel projection operator $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$, as described in Section II-D, is defined as
$\mathbf{P}(\hat{\mathbf{h}},\mathbf{a}):=\hat{\mathbf{h}}\hat{\mathbf{h}}^{\top}+\mathbf{a}\mathbf{a}^{\top}.$ (14)
According to (14), the vector $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$, used in (13), is the projection of the vector $\mathbf{x}_{\pi}$ onto the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$. Next, we discuss the design of the operator $\mathbf{R}$.
The rotation matrix $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a}):=R(\pi/2,\hat{\mathbf{h}},\mathbf{a})\in SO(n)$, where $R(\theta,\hat{\mathbf{h}},\mathbf{a})$ for $\theta\in[0,2\pi)$ is defined in Section II-E. For $\theta=\frac{\pi}{2}$, one gets
$\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})=\mathbf{I}_{n}+(\mathbf{a}\hat{\mathbf{h}}^{\top}-\hat{\mathbf{h}}\mathbf{a}^{\top})-(\hat{\mathbf{h}}\hat{\mathbf{h}}^{\top}+\mathbf{a}\mathbf{a}^{\top}).$ (15)
Note that, according to (14), one has $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, as per (15), the operator $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})$, used in (13), rotates the vector $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ by $\pi/2$ radians in the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$. This rotation is performed from the vector $\mathbf{h}$ to the vector $\mathbf{a}$.
Finally, the scalar function $\eta(\mathbf{x})\in[-1,1]$ is given by
$\eta(\mathbf{x})=\begin{cases}-1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\geq\gamma_{s},\\ 1-\frac{d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}-\gamma_{a}}{0.5(\gamma_{s}-\gamma_{a})},&\gamma_{a}<d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}<\gamma_{s},\\ 1,&d(\mathbf{x},\mathcal{O}_{\mathcal{W}})-r_{a}\leq\gamma_{a},\end{cases}$ (16)
where $0<\gamma_{a}<\gamma_{s}<\gamma$. The scalar function $\eta$ is designed to ensure that the center of the robot remains inside the $\gamma-$neighborhood of the $r_{a}-$dilated obstacle-occupied workspace $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$ when it operates in the obstacle-avoidance mode in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$. This feature allows for the design of the jump set of the obstacle-avoidance mode, as discussed later in Section IV-B2, and ensures convergence to the target location, as stated later in Theorem 1.
Remark 1.
Consider a plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ passing through the interior of obstacle $\mathcal{O}_{i}$ and the origin, where $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and $\mathbf{a}\in\mathbb{S}^{n-1}$. Since $\mathcal{O}_{i}$ is a general convex obstacle, one cannot guarantee that $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. Therefore, the operator $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$ is used in (13) to ensure that $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. The operator $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})$ aids in steering the robot along the boundary of obstacle $\mathcal{O}_{i}$ when it operates in the obstacle-avoidance mode. However, because $\mathcal{O}_{i}$ is a general convex obstacle, the vector $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ is not necessarily tangential to the set $\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$ at $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})$. Therefore, it may drive the robot either closer to obstacle $\mathcal{O}_{i}$ or away from it. Thus, the scalar function $\eta(\mathbf{x})$ in included to ensure that $\mathbf{x}$ remains in the set $\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ as long as the robot operates in obstacle-avoidance mode.
Next, we provide the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ used in (12).
IV-B Geometric construction of the flow and jump sets
When the robot is located at a distance larger than $\gamma$ from the obstacles, it moves straight towards the target location in the move-to-target mode. The robot’s distance from the nearby obstacles is obtained by evaluating $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})$, as discussed in Section II-B, where the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ is defined in (8). Upon entering in the $\gamma$-neighbourhood of the obstacles, as per Assumption 9, the robot is within the $\gamma$-neighbourhood of only one obstacle. In other words, according to Assumption 9, the fact that $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma$, implies the existence of $i\in\mathbb{I}$ such that $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and $\mathbf{x}\notin\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{j}))$ for all $j\in\mathbb{I}\setminus{i}$.
Figure 3: A workspace scenario in which condition (18) is satisfied.
Figure 4: Partitioning of the region $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, with $r_{a}=0.25\,m$ and $\gamma=1.15\,m$, into the avoidance region $\mathcal{R}_{a}^{i}$ and the exit region $\mathcal{R}_{e}^{i}$, for different values of the sensing radius $R_{s}$: (a) $R_{s}=0.5\,m$, (b) $R_{s}=1\,m$, (c) $R_{s}=1.5\,m$, (d) $R_{s}=4\,m$.
When the robot operates in the move-to-target mode, its velocity is directed towards the target location. Hence, if the robot enters in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$, it should constantly verify whether the path joining the robot’s location and the target is obstructed by $\mathcal{O}_{i}$. Therefore, we construct the set ð$\eth_{i}^{\mathbf{x}}$, which contains locations on the boundary of obstacle $\mathcal{O}_{i}$ that have a clear line of sight to the center of the robot. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, the set ð$\eth_{i}^{\mathbf{x}}$ is defined as
ð$\eth_{i}^{\mathbf{x}}:=\{\mathbf{y}\in\eth_{\mathcal{W}}^{\mathbf{x}}\,|\,\mathbf{y}\in\partial\mathcal{O}_{i}\}.$ (17)
Observe that if $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and
ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset,$ (18)
as shown in Fig. 3, then the robot, moving straight towards the target location, will eventually collide with obstacle $\mathcal{O}_{i}$, where $\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$ represents the cylinder with width $2r_{a}$ and axis $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})$ as defined in Section II-C5. We define the avoidance region, denoted by $\mathcal{R}_{a}^{i}$, as the set of all locations from the $\gamma-$neighbourhood of $r_{a}-$dilated obstacle $\mathcal{O}_{i}$ such that condition (18) is satisfied. The set $\mathcal{R}_{a}^{i}$ is defined as
$\mathcal{R}_{a}^{i}:=\left\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))|\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset\right\}.$ (19)
Since the set ð$\eth_{i}^{\mathbf{x}}$ is used to define the avoidance region $\mathcal{R}_{a}^{i}$, the shape of $\mathcal{R}_{a}^{i}$ depends on the value of the sensing radius $R_{s}$, as shown in Fig. 4. However, as stated in the next lemma, irrespective of the value of the sensing radius $R_{s}$, the avoidance region $\mathcal{R}_{a}^{i}$ is always a subset of the landing region $\mathcal{R}_{l}^{i}$, which is defined as
$\mathcal{R}_{l}^{i}:=\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))|\mathcal{L}_{s}(\mathbf{x},\mathbf{0})\cap$ $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\neq\emptyset,$ (20)
$\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0\}.$
Furthermore, if $R_{s}>l_{i}$, then one has $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$, where $l_{i}$ is the largest possible distance between any two points from the set $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, that is
$l_{i}=\max\{\lVert\mathbf{p}-\mathbf{q}\rVert|\mathbf{p},\mathbf{q}\in\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})\}.$ (21)
Lemma 1.
For each $i\in\mathbb{I}\setminus\{0\}$ and for any $R_{s}>0$, it holds that $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$. Additionally, if $R_{s}>l_{i}$, where $l_{i}$ is defined in (21), then $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$.
Proof.
See Appendix -A. ∎
In Fig. 4, the blue curve represents the boundary of the landing region $\mathcal{R}_{l}^{i}$ for obstacle $\mathcal{O}_{i}$. Notice that the avoidance region $\mathcal{R}_{a}^{i}$ is always a subset of the landing region $\mathcal{R}_{l}^{i}$, irrespective of the value of the sensing radius $R_{s}$.
Since $\mathcal{W}$ is a closed convex set, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$, condition (18) is not satisfied, where $\mathcal{O}_{0}=\left(\mathcal{W}^{\circ}\right)^{c}$. Therefore, $\mathcal{R}_{a}^{0}$ is an empty set. The union of the avoidance regions over all obstacles is given by
$\mathcal{R}_{a}:=\bigcup_{i\in\mathbb{I}}\mathcal{R}_{a}^{i}.$ (22)
Next, we define the exit region $\mathcal{R}_{e}$ as the part of the $\gamma-$neighbourhood of the $r_{a}-$dilated obstacles that do not belong the avoidance region. The exit region $\mathcal{R}_{e}$ is defined as
$\mathcal{R}_{e}=\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus\mathcal{R}_{a}.$ (23)
Note that when the robot’s center $\mathbf{x}$ belongs to the exit region, condition (18) is not satisfied. In other words, if the robot, with its center in the exit region, moves directly towards the target location along the path connecting its center to the origin, it will not collide with the nearest obstacle within the sensing radius $R_{s}$. Hence, the robot should move straight towards the target location only if it is in the exit region.
Next, we provide the geometric construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (12).
IV-B1 Flow and jump sets (move-to-target mode)
When the robot, in the move-to-target mode, enters the $\gamma$-neighborhood of an obstacle obstructing its straight path to the target location, the control input switches to the obstacle-avoidance mode. Hence, the jump set of the move-to-target mode for the state $\mathbf{x}$ is defined as
$\mathcal{J}_{0}^{\mathcal{W}}:={\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{R}_{a}},$ (24)
where $\gamma_{s}\in(0,\gamma).$ In (24), we allow the robot to enter the $\gamma_{s}$-neighborhood of obstacles before switching to the obstacle-avoidance mode. This creates a hysteresis region, $\mathcal{N}_{\gamma-\gamma_{s}}(\mathcal{D}_{r_{a}+\gamma_{s}}(\mathcal{O}_{\mathcal{W}}))$, which acts as a buffer zone to prevent frequent switching between the modes due to small disturbances or noise.
The flow set of the move-to-target mode for the state $\mathbf{x}$ is then defined as
$\mathcal{F}_{0}^{\mathcal{W}}:=\left(\mathcal{W}\setminus(\mathcal{D}_{r_{a}+\gamma_{s}}^{\circ}(\mathcal{O}_{\mathcal{W}}))\right)\cup\overline{\mathcal{R}_{e}}.$ (25)
Notice that the union of the jump set $\mathcal{J}_{0}^{\mathcal{W}}$ and the flow set $\mathcal{F}_{0}^{\mathcal{W}}$ covers the obstacle-free workspace $\mathcal{W}_{r_{a}}$, as shown in Fig. 5.
The flow set $\mathcal{F}_{0}$ and the jump set $\mathcal{J}_{0}$ for the move-to-target mode are given by
$\mathcal{F}_{0}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{F}_{0}^{\mathcal{W}},m=0\},$ (26)
$\mathcal{J}_{0}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{J}_{0}^{\mathcal{W}},m=0\}.$
Figure 5: Two-dimensional illustration of the flow sets and the jump sets considered in Sections IV-B1 and IV-B2.
IV-B2 Flow and jump sets (obstacle-avoidance mode)
In the obstacle-avoidance mode, the robot moves along the boundary of the obstacle being avoided in its $\gamma$-neighborhood. The control can switch back to the move-to-target mode only if the obstacle no longer obstructs the robot’s straight path to the origin. Thus, the switching occurs only when $\mathbf{x}$ is in the exit region $\mathcal{R}_{e}$ (23). Additionally, the hit point $\mathbf{h}$ (the robot’s position when it switched from the move-to-target to the current obstacle-avoidance mode) is used to ensure that, when the control input switches back to the move-to-target mode, the target location at the origin is closer to $\mathbf{x}$ than to $\mathbf{h}$. This condition helps us to establish the global convergence properties of the target location, as stated in Theorem 1. The jump set for the obstacle-avoidance mode $\mathcal{J}_{1}^{\mathcal{W}}$ is defined as follows:
$\mathcal{J}_{1}^{\mathcal{W}}:=\left(\mathcal{W}\setminus\mathcal{D}_{r_{a}+\gamma}^{\circ}(\mathcal{O}_{\mathcal{W}})\right)\cup\overline{\mathcal{ER}^{\mathbf{h}}}\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))\cup\mathcal{S}_{\mathbf{0}}.$ (27)
For any hit point $\mathbf{h}\in\mathcal{W}_{r_{a}}$, the set $\mathcal{ER}^{\mathbf{h}}$ is given by
$\mathcal{ER}^{\mathbf{h}}:=\{\mathbf{x}\in\mathcal{R}_{e}\big{|}\lVert\mathbf{h}\rVert-\lVert\mathbf{x}\rVert\geq\epsilon\},$ (28)
with $\epsilon\in(0,\bar{\epsilon}]$, where $\bar{\epsilon}$ is a sufficiently small positive scalar. The set $\mathcal{ER}^{\mathbf{h}}$ contains the locations $\mathbf{x}$ from the exit region $\mathcal{R}_{e}$ for which the target location is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$. Since the obstacles are compact and convex, and the target location $\mathbf{0}$ is within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}$, it is possible to guarantee the existence of the parameter $\bar{\epsilon}$ such that the intersection set $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ is non-empty for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for each $i\in\mathbb{I}\setminus\{0\}$, as stated in the following lemma:
Lemma 2.
Let Assumption 9 hold. Then, for each $i\in\mathbb{I}\setminus\{0\}$, for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, there exists $\bar{\epsilon}>0$ such that for any $\epsilon\in(0,\bar{\epsilon}]$, one has $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset,$ where $\mathcal{ER}^{\mathbf{h}}$ is defined in (28).
Proof.
See Appendix -B. ∎
According to (27) and (28), the control input corresponding to the obstacle-avoidance mode, can switch to the move-to-target mode when $\mathbf{x}$ belongs to the exit region $\mathcal{R}_{e}$ and the target location $\mathbf{0}$ is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$. This introduces a hysteresis region that prevents frequent switching between the modes due to small disturbances or noise.
Next, we define the set $\mathcal{S}_{0}$, used in (27).
Figure 6: (a) Illustration of the region $\mathcal{J}1^{\mathcal{W}}\setminus\mathcal{S}_{\mathbf{0}}$. (b) Illustration of the region $\mathcal{J}_{1}^{\mathcal{W}}$.
The set $\mathcal{S}_{\mathbf{0}}$, which contains locations from the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{B}_{R_{s}}(\mathbf{0})$ from which the robot can safely move straight towards the target location at the origin, is given by
$\mathcal{S}_{\mathbf{0}}:=\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{B}_{R_{s}}(\mathbf{0})|\eth_{\mathcal{W}}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset\}.$ (29)
Notice that if $\mathcal{S}_{\mathbf{0}}$ is excluded from the set $\mathcal{J}_{1}^{\mathcal{W}}$ in (27) and $\mathbf{0}$ is in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$ for some $i\in\mathbb{I}\setminus\{0\}$, then the $\mathbf{x}$ trajectory, starting from some $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, may not enter in the set $\overline{\mathcal{ER}^{\mathbf{h}}}$ and as such in the set $\mathcal{J}_{1}^{\mathcal{W}}$ when $\epsilon$ is set relatively high. This may cause the robot to indefinitely operate in the obstacle-avoidance mode in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$. For example, consider Fig. 6a, in which the $\mathbf{x}$ trajectory starting from $\mathbf{h}$ does not enter the set $\mathcal{J}_{1}^{\mathcal{W}}\setminus\mathcal{S}_{\mathbf{0}}$. Therefore, we include the set $\mathcal{S}_{\mathbf{0}}$ in (27) to ensure that the $\mathbf{x}$ trajectories starting from any $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ always enter in the set $\mathcal{J}_{1}^{\mathcal{W}}$ when the control input corresponds to the obstacle-avoidance mode, as shown in Fig. 6b.
The flow set of the obstacle-avoidance mode for the state $\mathbf{x}$, i.e., $\mathcal{F}_{1}^{\mathcal{W}}$, is defined as follows:
$\mathcal{F}_{1}^{\mathcal{W}}:=\overline{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus{(\mathcal{ER}^{\mathbf{h}}\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))\cup\mathcal{S}_{\mathbf{0}})}}.$ (30)
Notice that the union of the jump set (27) and the flow set (30) exactly covers the obstacle-free workspace $\mathcal{W}_{r_{a}}$, as shown in Fig. 5.
Remark 2.
Given that the workspace $\mathcal{W}$ is both convex and compact, there may exist some locations $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from which the nearest point to the robot’s center on the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ is not unique. This scenario prevents the implementation of the obstacle-avoidance term $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in the control law, at such locations. However, by excluding the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from the set $\mathcal{F}_{1}^{\mathcal{W}}$, as defined in (30), it is ensured that the obstacle-avoidance mode is never activated within the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$.
The flow set $\mathcal{F}_{1}$ and the jump set $\mathcal{J}_{1}$ for the obstacle-avoidance mode are given by
$\mathcal{F}_{1}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}},m=1,s\notin(s_{0},s_{0}+\delta_{s})\},$ (31)
$\mathcal{J}_{1}:=\mathcal{J}_{\mathbf{x}}\cup\mathcal{J}_{s},$
where $\mathcal{J}_{\mathbf{x}}:=\{\xi\in\mathcal{K}|m=1,\mathbf{x}\in\mathcal{J}_{1}^{\mathcal{W}}\}$ and $\mathcal{J}_{s}:=\{\xi\in\mathcal{K}|m=1,s\in[s_{0},s_{0}+\delta_{s}]\}$ with $s_{0}\in\mathbb{R}_{\geq 0}$ being the initial value of the state $s$, i.e., $s(0,0)=s_{0}$, and $0<\delta_{s}<\tau_{s}$ for some $\tau_{s}>0$.
Remark 3.
The definition of sets in (31) enables the control to immediately switch to the move-to-target mode if it is initialized in the obstacle-avoidance mode (i.e., $\xi(0,0)\in\mathcal{J}_{1}$). This ensures that the hit point $\mathbf{h}$ always belongs to the set $\mathcal{J}_{0}^{\mathcal{W}}$ before the robot starts moving in the obstacle-avoidance mode, thereby confirming the existence of the parameter $\bar{\epsilon}$, as established in Lemma 2.
Finally, the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (12), are defined as
$\mathcal{F}:=\bigcup_{m\in\mathbb{M}}\mathcal{F}_{m},\quad\mathcal{J}:=\bigcup_{m\in\mathbb{M}}\mathcal{J}_{m},$ (32)
where $\mathcal{F}_{0},\mathcal{J}_{0}$ defined in (26) for $m=0$ and in (31) for $m=1.$ Next, we provide the update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ used in (12b).
IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$
The update law $\mathbf{L}(\xi)$, used in (12b), updates the value of the hit point $\mathbf{h}$, the unit vector $\mathbf{a}$, the mode indicator $m$ and the variable $s$ when the state $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$ belongs to the jump set $\mathcal{J}$ defined in (32) and is given by
$\mathbf{L}(\xi)=\begin{cases}\mathbf{L}_{0}(\xi),&\xi\in\mathcal{J}_{0},\\ \mathbf{L}_{1}(\xi),&\xi\in\mathcal{J}_{1}.\end{cases}$ (33)
When the state $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, defined in (26), the update law $\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$ is given as
$\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{a}^{\prime}\\ 1\\ s+\tau_{s}\end{bmatrix},\mathbf{a}^{\prime}\in\mathbf{A}(\mathbf{x})\right\},$ (34)
where $\tau_{s}>0$. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x})\cap\mathcal{P}(\mathbf{x},\mathbf{x}_{\pi}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (35)
where for any $\mathbf{p}\in\mathbb{R}^{n}$, the set $\mathcal{P}^{\perp}(\mathbf{p})$, which is defined as
$\mathcal{P}^{\perp}(\mathbf{p}):=\{\mathbf{q}\in\mathbb{S}^{n-1}|\mathbf{q}^{\intercal}\mathbf{p}=0\},$ (36)
contains unit vectors that are perpendicular to the vector $\mathbf{p}.$
As per (34), when $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, the current value of $\mathbf{x}$ is assigned to the hit point $\mathbf{h}$. Moreover, using (35), the unit vector $\mathbf{a}$ is updated to be perpendicular to $\mathbf{x}$ and lies in the plane spanned by $\mathbf{x}$ and $\mathbf{x}_{\pi}$, provided they are not collinear. As a result, the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ passes through the target location at the origin and intersects the interior of the obstacle being avoided. This ensures that $\mathcal{P}(\mathbf{h},\mathbf{a})$ intersects both the avoidance region $\mathcal{R}_{a}$ and the exit region $\mathcal{R}_{e}$ associated with the obstacle to be avoided. This property guarantees that, while operating in the obstacle-avoidance mode, the robot eventually enters in the exit region where the obstacle no longer blocks its straight path to the origin, and switches back to the move-to-target mode. This allows one to establish global convergence properties of the target location at the origin, as stated later in Theorem 1.
When the state $\xi$ enters in the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode, defined in (31), the update law $\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)$ is given by
$\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)=\begin{bmatrix}\mathbf{h}\\ \mathbf{a}\\ 0\\ s+\tau_{s}\end{bmatrix}.$ (37)
According to (37), when the robot switches from the obstacle-avoidance mode to the move-to-target mode, the coordinates of the hit point $\mathbf{h}$ and the unit vector $\mathbf{a}$ remain unchanged.
Remark 4.
Since the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 9, the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint. Furthermore, according to (30), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
Control design summary: The proposed hybrid feedback control law can be summarized as follows:
  • Parameters selection: the target location is set at the origin with $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$, and $\xi(0,0)\in\mathcal{K}$. The gain parameters $\kappa_{s}$ and $\kappa_{r}$ are set to positive values, and a sufficiently small positive value is chosen for $\bar{\epsilon}$, used in (28). The scalar parameter $\gamma$, used in the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, is selected such that $\gamma\in(0,\delta-r_{s})$. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ The parameter $\tau_{s}$ is selected such that $\tau_{s}>0$ and $\delta_{s}$, used in (31), is set to ensure that $0<\delta_{s}<\tau_{s}$.
  • Move-to-target mode $m=0$: this mode is activated when $\xi\in\mathcal{F}_{0}$. As per (12a), the control input is given by $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x}$, causing $\mathbf{x}$ to evolve along the line segment $\mathcal{L}_{s}(\mathbf{0},\mathbf{x})$ towards the origin. If, at some instance of time, $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode, the state variables $(\mathbf{h},\mathbf{a},m,s)$ are updated using (34), and the control input switches to the obstacle-avoidance mode.
  • Obstacle-avoidance mode $m=1$: this mode is activated when $\xi\in\mathcal{F}_{1}$. As per (12a), the control input is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, causing $\mathbf{x}$ to evolve in the $\gamma-$neighborhood of the nearest obstacle along the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ until the state $\xi$ enters in the jump set $\mathcal{J}_{1}$ of the obstacle-avoidance mode. When $\xi\in\mathcal{J}_{1}$, the state variables $(m,s)$ are updated using (37), and the control input switches to the move-to-target mode.
This concludes the design of the proposed hybrid feedback controller (12). Next, we analyze the safety, stability and convergence properties of the proposed hybrid feedback controller.
V Stability Analysis
The hybrid closed-loop system resulting from the hybrid feedback control law (12) is given by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}(\xi),{\xi\in\mathcal{F}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}(\xi),{\xi\in\mathcal{J}}},$ (38)
where $\mathbf{u}(\xi)$ is defined in (12a), and the update law $\mathbf{L}(\xi)$ is provided in (33). Definitions of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ are provided in (32). Next, we analyze the hybrid closed-loop system (38) in terms of forward invariance of the obstacle-free state space $\mathcal{K}$, along with the stability properties of the target set
$\mathcal{A}:=\{\xi\in\mathcal{K}|\mathbf{x}=\mathbf{0}\}.$ (39)
The next lemma shows that the hybrid closed-loop system (38) satisfies the hybrid basic conditions [27, Assumption 6.5], which guarantees the well-posedness of the hybrid closed-loop system.
Lemma 3.
The hybrid closed-loop system (38) with data $(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J})$ satisfies the following hybrid basic conditions:
  1. 1.
    the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (32), are closed subsets of $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}$.
  2. 2.
    the flow map $\mathbf{F}$, defined in (38), is outer semicontinuous and locally bounded relative to $\mathcal{F}$, dom $\mathcal{F}\subset\text{dom }\mathbf{F},$ and $\mathbf{F}(\xi)$ is convex for every $\xi\in\mathcal{F},$
  3. 3.
[…]
Proof.
See Appendix -C. ∎
For safe autonomous navigation, the state $\mathbf{x}$ must always evolve within the obstacle-free workspace $\mathcal{W}_{r_{a}}$, defined in (10). This is equivalent to having the set $\mathcal{K}$ forward invariant for the hybrid closed-loop system (38). This is stated in the next Lemma.
Lemma 4.
Under Assumption 9, for the hybrid closed-loop system (38), the obstacle-free set $\mathcal{K}:=\mathcal{W}_{r_{a}}\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ is forward invariant.
Proof.
See Appendix -D. ∎
Next, we provide one of our main results which states that, for all initial conditions in the obstacle-free set $\mathcal{K}$, the proposed hybrid controller not only ensures safe navigation but also guarantees global asymptotic stability of the target location at the origin.
Theorem 1.
Under Assumption 9, for the hybrid closed-loop system (38), the following holds true:
  1. i)
    the obstacle-free set $\mathcal{K}$ is forward invariant;
  2. ii)
    the target set $\mathcal{A}$ is globally asymptotically stable over the set $\mathcal{K}$;
  3. iii)
    the number of jumps is finite.
Proof.
See Appendix -E. ∎
VI Application to sphere worlds
Obviously, the hybrid feedback controller (12), which is designed for safe autonomous navigation in $n-$D environments with arbitrary convex obstacles, is applicable in sphere words i.e., environments with spherical obstacles. However, in this section, we will take advantage of the simplified geometry of the obstacles to redesign the control law (12) in a way that ensures a monotonic decrease of the distance between the robot and the target location–a feature that is not guaranteed with the control law (12) in environments with arbitrary convex obstacles.
Let us consider environments with spherical obstacles as defined in [17]. The workspace $\mathcal{W}:=\mathcal{B}_{r_{0}}(\mathbf{c}_{0})$ is a compact sphere with radius $r_{0}\in\mathbb{R}_{>0}$ and center $\mathbf{c}_{0}\in\mathbb{R}^{n}.$ In addition, the workspace $\mathcal{W}$ contains disjoint, compact spherical obstacles $\mathcal{O}_{i}:=\mathcal{B}_{r_{i}}(\mathbf{c}_{i}),i\in\mathbb{I}\setminus\{0\}$, where $r_{i}\in\mathbb{R}_{\geq 0}$ and $\mathbf{c}_{i}\in\mathcal{W}$ represent the radius and the center of obstacle $\mathcal{O}_{i}.$ Similar to [17], the workspace $\mathcal{W}$ satisfies Assumption 9.
Taking advantage of the spherical geometry of the obstacles, we will appropriately design the unit vector $\mathbf{a}$ and modify the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in (13), to ensure that in the obstacle-avoidance mode, the distance between the target location and the robot is monotonically decreasing.
Figure 7: The left figure depicts the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{P}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, shaded in a dark blue color. The right figure shows the direction of the modified obstacle-avoidance control vector, $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, at $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{P}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$.
For any $\mathbf{p}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, we define a set-valued mapping $\mathbf{A}_{s}$ as follows:
$\mathbf{A}_{s}(\mathbf{p})=\{\mathbf{q}\in\mathbf{A}(\mathbf{p})|\mathbf{q}^{\top}\mathbf{p}_{\pi}\geq 0\},$ (40)
where the set-valued mapping $\mathbf{A}$ is defined in (35), and $\mathbf{p}_{\pi}=\mathbf{p}-\Pi(\mathbf{p},\mathcal{O}_{\mathcal{W}})$.
Note that for any hit point $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for $i\in\mathbb{I}$, if one sets $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$, then, since obstacle $\mathcal{O}_{i}$ is a sphere, it can be shown that for all $\mathbf{x}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ $\Pi(\mathbf{x},\mathcal{O}_{i})$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Therefore, for any given $\mathbf{h}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$, $\beta\in[r_{a},r_{a}+\gamma]$ and for all $\mathbf{x}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one can show that the vector $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ belongs to the intersection of the hyperplane $\mathcal{H}(\mathbf{0},\mathbf{x}_{\pi})$ and the plane $\mathcal{P}(\mathbf{h},\mathbf{a}),$ where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$. This allows one to ensure that if $\mathbf{h}\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})$ for $\beta\in[r_{a},r_{a}+\gamma]$, and $\mathbf{x}(t_{0},j_{0})\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for some $(t_{0},j_{0})\in\text{ dom }\xi$, where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$, then under the control input $\mathbf{u}(\xi)=\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, where
$\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})=\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},$ (41)
one has $\mathbf{x}(t,j)\in\partial\mathcal{D}_{\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all time $(t,j)\succeq(t_{0},j_{0}),$ ensuring robot’s safety, as stated later in Theorem 2
Now, for a given $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, consider the hyperplane $\mathcal{H}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$ and $\mathbf{c}_{i}$ is the center of obstacle $\mathcal{O}_{i}$. Notice that, for all locations $\mathbf{x}$ in the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, which is depicted in Fig. 7, the inner product between the modified obstacle-avoidance control term $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ and the vector $\mathbf{x}$ is always non-positive. This allows one to show that if $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$ and $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, then the control input vector $\mathbf{u}(\xi)=\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ will ensure monotonic decrease of the distance $\|{\mathbf{x}}\|$ as long as the state $\mathbf{x}$ remains in the set $\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i})$, as stated later in Theorem 2.
Next, motivated by the preceding discussion, we modify the proposed hybrid feedback control law (12) using the modified obstacle-avoidance control vector $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, defined in (41), and the set-valued mapping $\mathbf{A}_{s}$, provided in (40), as follows:
  1. 1.
    the modified hybrid control input vector $\mathbf{u}_{s}(\xi)$, which is obtained by replacing the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ with the modified obstacle-avoidance control vector $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ in (12a), is given as
    $\mathbf{u}_{s}(\xi)=-\kappa_{s}(1-m)\mathbf{x}+\kappa_{r}m\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a}),$ (42)
    where $\kappa_{s}>0$ and $\kappa_{r}>0.$
  2. 2.
    the modified update law $\mathbf{L}_{0}^{s}(\xi)$ when the state $\xi$ enters in the jump set $\mathcal{J}_{0}$ of the move-to-target mode is given as
    $\mathbf{L}_{0}^{s}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{a}^{\prime}\\ 1\\ s+\tau_{s}\end{bmatrix},\mathbf{a}^{\prime}\in\mathbf{A}_{s}(\mathbf{x})\right\},$ (43)
    where the set-valued mapping $\mathbf{A}_{s}$ is defined according to (40).
The hybrid closed-loop system resulting from the modified hybrid feedback control law is given by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}_{s}(\xi)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\end{matrix}}_{\dot{\xi}=\mathbf{F}_{s}(\xi),{\xi\in\mathcal{F}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}_{s}(\xi)\end{matrix}}_{\xi^{+}\in\mathbf{J}_{s}(\xi),{\xi\in\mathcal{J}}},$ (44)
where the control input vector $\mathbf{u}_{s}$ is defined in (42) and the update law $\mathbf{L}_{s}$ is obtained by replacing $\mathbf{L}_{0}$ in (33) with $\mathbf{L}_{0}^{s}$ given in (43).
The next lemma shows that the hybrid closed-loop system (44) with the data $(\mathcal{F},\mathbf{F}_{s},\mathcal{J},\mathbf{J}_{s})$ satisfies the hybrid basic conditions, as stated in Lemma 3.
Lemma 5.
The hybrid closed-loop system (44) with data $(\mathcal{F},\mathbf{F}_{s},\mathcal{J},\mathbf{J}_{s})$ satisfies the hybrid basic conditions stated in Lemma 3.
The proof of Lemma 5 is similar to the proof of Lemma 3, therefore, it is omitted.
Next, we demonstrate that for the robot operating in a sphere world, which satisfies Assumption 9, the modified proposed hybrid feedback controller ensures safe navigation. It also guarantees global asymptotic stability of the target location at the origin, with a monotonic decrease in the distance between the robot’s center and the target location.
Theorem 2.
Under Assumption 9, for the hybrid closed-loop system (44), the following holds true:
  1. i)
    the obstacle-free set $\mathcal{K}$ is forward invariant,
  2. ii)
    the target set $\mathcal{A}$ is globally asymptotically stable over the set $\mathcal{K}$,
  3. iii)
    the number of jumps is finite,
  4. iv)
    the distance from the robot’s location to the target location is monotonically decreasing.
Proof.
See Appendix -G. ∎
VII Implementation Procedure
We consider a workspace with convex obstacles that satisfies Assumption 9 with some $\delta>0$, as discussed in Section III. The target location is set at the origin within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$. The parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ are chosen to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma<(\delta-r_{s})$. A sufficiently small value for $\bar{\epsilon}$ is selected, and the parameter $\epsilon$, used in (28), is chosen such that $\epsilon\in(0,\bar{\epsilon}]$. The parameter $\delta_{s}$, used in (31), is set such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$. The robot is equipped with a range scanner with a sensing radius of $R_{s}>r_{a}+\gamma$. The composite state vector $\xi$ is initialized in the set $\mathcal{K}$.
Switching from the move-to-target mode to the obstacle-avoidance mode: When the control input is initialized in the move-to-target mode, according to (12a), it steers the robot straight towards the origin. The robot should constantly measure the distance between its center and the surrounding obstacles to identify whether $\xi$ has entered in the jump set $\mathcal{J}_{0}$ of the move-to-target mode. To do this, the robot needs to identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$, as defined in (8), which contains the locations from the boundary of the surrounding obstacles that are less than $R_{s}$ units away from the center of the robot and have a clear line of sight to the center of the robot, where $R_{s}>r_{a}+\gamma$ represents the sensing radius. Then, one can obtain the distance between the robot’s center and the surrounding obstacles by evaluating $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})$ according to Section II-B.
If $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma_{s}$, one should identify whether the robot can move straight towards the target location without colliding with the nearest obstacle within the sensing region. In other words, one should check condition (18) to identify whether the center of the robot belongs to the avoidance region $\mathcal{R}_{a}$, defined in (19), associated with the nearest obstacle, let us say $\mathcal{O}_{i},i\in\mathbb{I}$. To that end, one needs to identify the set ð$\eth_{i}^{\mathbf{x}}$, as defined in (17), which contains the locations from the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ that belong to the boundary of the closest obstacle $\mathcal{O}_{i}$. Once the set ð$\eth_{i}^{\mathbf{x}}$ has been identified, one needs to determine whether $\mathbf{x}$ belongs to the avoidance region $\mathcal{R}_{a}$, defined in (19), by evaluating the intersection between the set ð$\eth_{i}^{\mathbf{x}}$ and the set $\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$. If ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\not=\emptyset$, then the center of the robot belongs to the avoidance region $\mathcal{R}_{a}$ and the state $\xi$ has entered in the jump set $\mathcal{J}_{0}$ of the move-to-target. Otherwise, the robot continues to operate in the move-to-target mode.
Switching from the obstacle-avoidance mode to the move-to-target mode: When the state $\xi$ enters in the jump set $\mathcal{J}_{0}$, the state $\xi$ is updated as per (34) and (38), and the control input switches to the obstacle-avoidance mode. According to Lemma 6, when the robot operates in the obstacle-avoidance mode, it stays inside the $\gamma-$neighborhood of the closest obstacle. As the robot operates in the obstacle-avoidance mode, we continuously evaluate the intersection ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$ to check whether the center of the robot has entered in the exit region $\mathcal{R}_{e}$. If ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset$, then the center of the robot belongs to the exit region $\mathcal{R}_{e}$. Additionally, if the target location is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$, it implies that $\xi\in\mathcal{J}_{1}$. Then, the value of the variables $m$ and $s$ are updated as per (37) and the control input switches to the move-to-target mode.
Finally, if the control input is initialized in the obstacle-avoidance mode, according to (31), the state $\xi(0,0)$ belongs to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$. As a result, according to (37), the variables $\mathbf{h},\mathbf{a},m$ and $s$, are updated and the control input switches to the move-to-target mode.
The above-mentioned implementation procedure is summarized in Algorithm 1.
VII-A Control-input smoothing mechanism for practical implementations
Since the proposed hybrid closed-loop system (38) satisfies the hybrid basic conditions, as mentioned in Lemma 3, the resultant robot trajectories during the flow are always smooth. However, the control input vector $\mathbf{u}$, defined in (12a), in general, changes direction instantaneously when switching between modes, causing the overall robot trajectories to become non-smooth. This instantaneous change in the direction of $\mathbf{u}$ is caused by a discrete change in the value of the mode indicator variable $m$ when the state $\xi$ enters in the jump set $\mathcal{J}$. Therefore, in order to generate smooth robot trajectories, one solution consists in replacing $m$ in (12a) with the following continuous scalar function $\lambda$ which takes values between 00 and $1$:
$\lambda(m,s,\tau)=\begin{cases}\varphi_{t_{s}}(s-\tau),&m=1,\\ \varphi_{t_{s}}(\tau-s),&m=0,\end{cases}$ (45)
where the scalar variable $\tau$ is kept constant during the flow and is updated only when $\xi$ enters in the jump set $\mathcal{J}$. The update law for $\tau$, which is denoted by $\mathbf{U}(\xi)$ is given by
$\mathbf{U}(\xi)=\begin{cases}s+\tau_{s},&\xi\in\mathcal{J}_{sm}^{0},\\ s+\tau_{s}+t_{s},&\xi\in\mathcal{J}_{sm}^{1},\end{cases}$ (46)
with $t_{s}>0$ and $\tau_{s}>0$ as defined in (34). For $z\in\{0,1\}$, the jump set $\mathcal{J}_{sm}^{z}$ is defined as
$\mathcal{J}_{sm}^{z}:=\{\xi\in\mathcal{J}_{z}\mid s-\tau\geq t_{s}\},$ (47)
where $\mathcal{J}_{z}$ for $z=0$ and $z=1$ is given by (26) and (31), respectively. According to (46) and (47), $\tau$ is updated only when $\xi$ enters in the jump set $\mathcal{J}_{sm}$ and the difference $s-\tau$ is greater than or equal to $t_{s}$ seconds, where $\mathcal{J}_{sm}=\mathcal{J}_{sm}^{0}\cup\mathcal{J}_{sm}^{1}$. This ensures that the consecutive switching instances are separated by at least $t_{s}$ seconds. According to (46) and (47), the variable $\tau$ is updated only when $\xi$ enters the jump set $\mathcal{J}$ and the condition $s-\tau\geq t_{s}$ is satisfied. This ensures that consecutive switching events are separated by at least $t_{s}$ seconds.
Given a scalar variable $p$, the continuous scalar function $\varphi_{t_{s}}(p)$ is defined by
$\varphi_{t_{s}}(p)=\begin{cases}1,&p\geq t_{s},\\ p/t_{s},&0\leq p\leq t_{s},\\ 0,&p\leq 0.\end{cases}$ (48)
Depending on the value of $p$, the function $\varphi_{t_{s}}(p)$ takes values between 00 and $1$.
Notice that according to (34) and (46), when $\xi\in\mathcal{J}_{0}$, the variables $s$ and $\tau$ are updated to $s+\tau_{s}$. Since $\tau$ is kept constant during the flow, one has $\dot{\tau}=0$. Additionally, as per (38), we have $\dot{s}=1$. Therefore, as the robot enters in the obstacle-avoidance mode, the difference $s-\tau$ equals zero and keeps increasing until the robot switches to the move-to-target mode. This, according to (45) and (48), causes the value of $\lambda$ to go from 00 to $1$ in $t_{s}$ seconds [^2] The parameter $t_{s}$ should be set to a relatively small positive value to ensure that, depending on the current mode of operation, the function $\lambda$ reaches a constant value of either 00 or $1$ before the state $(\xi,\tau)$ enters the jump set of the respective mode.. Similarly, according to (34) and (46), when $\xi\in\mathcal{J}_{1}$, the variables $s$ and $\tau$ are updated to $s+\tau_{s}$ and $s+\tau_{s}+t_{s}$, respectively. Addtionally, during the flow we have $\dot{s}=1$ and $\dot{\tau}=0$. Therefore, as the robot enters in the move-to-target mode, the difference $\tau-s$ equals $t_{s}$ and keeps decreasing until the robot switches to the obstacle-avoidance mode. This, according to (45) and (48), causes the value of $\lambda$ to go from $1$ to 00 in $t_{s}$ seconds. Therefore, by replacing $m$ in (12a) with $\lambda(m,s,\tau)$ one can ensure that the control input vector changes its value between $-\kappa_{s}\mathbf{x}$ and $\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in a continuous manner, depending on the current mode of operation. The smoothed version of the hybrid control input is then given by
$\mathbf{u}_{sm}(\xi,\tau)=-\kappa_{s}(1-\lambda(m,s,\tau))\mathbf{x}+\kappa_{r}\lambda(m,s,\tau)\mathbf{v}.$ (49)
The resulting hybrid closed-loop system is represented by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\\ \dot{\tau}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}_{sm}(\xi,\tau)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\\ \;=0\end{matrix}}_{(\dot{\xi},\dot{\tau})=\mathbf{F}_{sm}(\xi,\tau),{(\xi,\tau)\in\mathcal{F}\times\mathbb{R}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\\ \tau^{+}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\\ =\mathbf{U}(\xi)\end{matrix}}_{(\xi^{+},\tau^{+})\in\mathbf{J}_{sm}(\xi),{(\xi,\tau)\in\mathcal{J}_{sm}\times\mathbb{R}}}$ (50)
where $\mathcal{J}_{sm}=\mathcal{J}_{sm}^{0}\cup\mathcal{J}_{sm}^{1}$. Notice that the variable $\tau$ does not change value during the flow, and is updated according to $\mathbf{U}(\xi)$ (46) only when $\xi\in\mathcal{J}_{sm}$.
Remark 5.
The control input vector $\mathbf{u}_{sm}$ allows one to generate smooth robot trajectories that converge to the target location at the origin as long as the robot’s location $\mathbf{x}$ and the mode indicator $m$ are not initialized as follows:
  • $\mathbf{x}(0,0)\in\partial\mathcal{W}_{r_{a}}\cap\mathcal{F}_{1}^{\mathcal{W}}$ and $m(0,0)=0$. In this case, one cannot ensure the forward invariance of the set $\mathcal{K}\times\mathbb{R}$ as $\mathbf{u}_{sm}(\xi(0,0),\tau(0,0))$ at $\mathbf{x}(0,0)$ is directed towards the interior of set $\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})$.
  • $\mathbf{x}(0,0)\in\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$ and $m(0,0)=1$. In this case, the obstacle-avoidance control input vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ is not viable as the uniqueness of $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is not guaranteed.
To avoid such situations, one should initialize the robot’s location $\mathbf{x}$ in the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}$ and ensure that the robot starts operating in the move-to-target mode with the stabilization control input vector $-\kappa_{s}\mathbf{x}$. This can be achieved by setting $m(0,0)=0$ and choosing $\tau(0,0)\in(-\infty,s(0,0))$, as per (45) and (49).
Algorithm 1 Implementation of the proposed hybrid control law (12) in a priori unknown environment.
1:  Set target location at the origin $\mathbf{0}.$
2:  Initialize $\mathbf{x}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{h}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{a}(0,0)\in\mathbb{S}^{n-1}$, $m(0,0)\in\mathbb{M}$ and $s(0,0)\in\mathbb{R}_{\geq 0}$. Choose a sufficiently small value of $\bar{\epsilon}$ according to Lemma 2, and initialize $\epsilon\in(0,\bar{\epsilon}].$ Select a minimum safety distance $r_{s}$ such that $r_{s}\in(0,\delta)$ and set the parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ such that $0<\gamma_{a}<\gamma_{s}<\gamma<\delta-r_{s}$. Choose $R_{s}>r_{a}+\gamma$, used in (8). The parameter $\delta_{s}$, used in (31), is set such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$.
3:  Measure $\mathbf{x}$ and the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (8).
4:  if $m=0$then
5:     if $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma_{s}$then
6:        Identify the set ð$\eth_{i}^{\mathbf{x}}\subset\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (17).
7:        if ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset$then
8:           Update $\xi\leftarrow\mathbf{J}(\xi)$ using (34) and (38).
9:        end if
10:     end if
11:  end if
12:  if $m=1$then
13:     if $s=s(0,0)$then
14:        Update $\xi\leftarrow\mathbf{J}(\xi)$ using (37) and (38).
15:     else
16:        if $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+{\gamma},$ then
17:           Identify the set ð$\eth_{i}^{\mathbf{x}}\subset\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (17).
18:           if ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset$then
19:              if $\lVert\mathbf{x}\rVert\leq\lVert\mathbf{h}\rVert-\epsilon,$ then
20:                 Update $\xi\leftarrow\mathbf{J}(\xi)$ using (37) and (38).
21:              end if
22:           end if
23:        else
24:           Update $\xi\leftarrow\mathbf{J}(\xi)$ using (37) and (38).
25:        end if
26:     end if
27:  end if
28:  Execute $\mathbf{F}(\xi)$ given (12), used in (38).
29:  Go to step 3.
VIII Simulation Results
We consider an unbounded workspace i.e., obstacle $\mathcal{O}_{0}=\emptyset$, with $2$ three-dimensional, convex obstacles, as shown in Fig. 8. We apply the proposed hybrid feedback controller (12) for the robot initialized at $8$ different locations in the obstacle-free workspace. The target is located at the origin. The radius of the robot is set to $0.2m$. The minimum safety distance $r_{s}=0.1m$ and the parameter $\gamma=0.5m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $1$ and $2$, respectively. The sensing radius $R_{s}$, used in (8), is set to $2m$. The parameter $\epsilon$, used in (28), is set to $0.5m$. From Fig. 8, it can be observed that the robot converges to the target location while simultaneously avoiding collisions with the obstacles. Fig. 9 shows that the center of the robot stays at least $r_{a}$ meters away from the boundary of the obstacles.
Figure 8: Robot trajectories starting from different locations.
Figure 9: Distance of the center of the robot $\mathbf{x}$ from the boundary of the obstacle occupied workspace $\mathcal{O}_{\mathcal{W}}$.
We now consider a two-dimensional workspace containing 12 convex obstacles. A point robot is initialized at 24 different locations along the workspace boundary, as depicted in Fig. 10, with the target located at the origin. The safety distance $r_{s}$ is set to $0.1\,m$. The parameters $\gamma$ and $\epsilon$ are set to $0.2\,m$ and $0.5\,m$, respectively. The gain parameters are $\kappa_{s}=1$ and $\kappa_{r}=2$, and the sensing radius $R_{s}$ is set to $1\,m$. In Fig. 10, dashed trajectories represent the robot’s motion in the move-to-target mode, while solid trajectories indicate motion in the obstacle-avoidance mode. Under the hybrid feedback controller (12), the robot safely converges to the target from all initial locations.
Figure 10: Safe convergence of robot trajectories to the target location.
Next, we examine the effect of varying the sensing radius $R_{s}$ on the robot’s trajectories. The robot with radius $r=0.15m$ is initialized at $[0,9.5]^{\top}$ in a two-dimensional workspace with a single convex obstacle, as shown in Fig. 11. The target location is set at the origin. The safety distance $r_{s}$ is set to $0.05m$. The parameters $\gamma$ and $\epsilon$ are set to $1.5m$ and $0.5m$, respectively. The gains are $\kappa_{s}=1$ and $\kappa_{r}=2$. In Fig. 11, the blue portions of the trajectories correspond to the motion of the robot in move-to-target mode, while red portions correspond to the motion in the obstacle-avoidance mode. When the sensing radius is relatively small, the robot cannot detect all points on the boundary of obstacle $\mathcal{O}_{i}$ that are within its line of sight but outside the sensing region $\mathcal{B}_{R_{s}}(\mathbf{x})$. As a result, the robot has only partial information about the obstacle’s boundary. Consequently, the robot may switch to the move-to-target mode even when it does not have a clear line of sight to the target, as shown in Figs. 11a, 11b, and 11c.
Figure 11: Robot trajectories for different values of the sensing radius $R_{s}$: (a) $R_{s}=1.7m$, (b) $R_{s}=2.5m$, (c) $R_{s}=3.4m$, (d) $R_{s}=4m$.
We next provide a comparison with the separating hyperplane approach developed in [15]. Similar to our approach, this approach can be implemented in a priori unknown environments using the information obtained via a range-bearing sensor mounted on the robot. Contrary to our approach, this approach only works for convex obstacles that satisfy the obstacle curvature condition [15, Assumption 2]. When the workspace consists of obstacles that do not satisfy the obstacle curvature condition, the separating hyperplane approach generates trajectories that converge to an undesired equilibrium (local minimum), as shown in Fig. 12b. On the other hand, the proposed hybrid feedback controller guarantees safe, global asymptotic convergence to the target location, as seen in Fig. 12a. The complete simulation video can be found at333[Online]. Available: https://youtu.be/R6OowaF6dTc.
Figure 12: (a) Robot trajectories obtained using our proposed hybrid feedback approach. (b) Robot trajectories obtained using the separating hyperplane approach [15].
Figure 13: Robot trajectories safely navigating around spherical obstacles and converging to the target location at the origin.
In the next simulation scenario, as shown in Fig. 13, we consider 2 three-dimensional spherical obstacles and apply the proposed hybrid controller (12) with modifications mentioned in Section VI, for a point robot initialized at $8$ different locations in the obstacle-free workspace. The safety distance $r_{s}$ is set to $0.15m$ and the parameter $\gamma=0.15m.$ The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $0.5$. The sensing radius $R_{s}$, used in (8), is set to $2m$. The parameter $\epsilon$, used in (28), is set to $0.05m$. Similar to the previous simulations, it can be observed from Fig. 14 that the robot converges asymptotically to the target location without colliding with the obstacles. Since the obstacles are spheres, the distance between the robot and the target location monotonically decreases as the robot converges to the target location, as stated in Theorem 2 and as shown in Fig. 14b.
Figure 14: (a) Distance between the robot $\mathbf{x}$ and the obstacles $\mathcal{O}_{\mathcal{W}}$ versus time. (b) Distance between the robot $\mathbf{x}$ and the target location versus time.
Figure 15: (a) Robot trajectory for the hybrid closed-loop system (38) using the control input vector $\mathbf{u}(\xi)$. (b) Robot trajectory for the smoothed hybrid closed-loop system (50) using the control input vector $\mathbf{u}_{sm}(\xi,\tau)$.
We compare the robot trajectories generated by the hybrid closed-loop system (38) with the smoothed robot trajectories obtained from the modified hybrid closed-loop system (50). The environment contains two convex obstacles and the target location is positioned at the origin, as shown in Fig. 15. The robot, with a radius of $r=0.175m$, is initialized at the point $[5.5,3]^{\top}$. The safety distance $r_{s}$ is set to $0.025m$, and the parameter $\gamma$ is set to $0.3m$. The control gains $\kappa_{s}$ and $\kappa_{r}$ are set to 0.2 and 2, respectively. The sensing radius $R_{s}$ is set to $3m$, and the parameter $\epsilon$ is set to $1m$. The time parameters $t_{s}$ and $\tau_{s}$, used in (46), are set to $0.5$ seconds and $1$ second, respectively.
In Fig. 15, the red portions of the trajectories represent the motion of the robot in the move-to-target mode, whereas the blue portions correspond to the motion in the obstacle-avoidance mode. Figure 16 shows the time-evolution of the magnitude of the control input, while Fig. 17 depicts the direction of the control input represented by the angle between the control input vector and the $x$-axis. Given $\mathbf{p}=[p_{1},p_{2}]^{\top}\in\mathbb{R}^{2}$, the angle between $\mathbf{p}$ and the $x$-axis is computed as:
$\mathbf{atan2v}(\mathbf{p})=\mathbf{mod}(\mathbf{atan2}(p_{2}/p_{1})+2\pi,2\pi).$ (51)
In Fig. 16 and Fig. 17, the red trajectories represent the control input vector $\mathbf{u}$, while the blue trajectories correspond to $\mathbf{u}_{sm}$.
Generally, the magnitude and direction of the control input vector $\mathbf{u}$ changes instantaneously when switching between modes, as illustrated in Fig. 16 and Fig. 17, respectively. This can sometimes cause an abrupt change in the robot’s direction of motion when the control input $\mathbf{u}$ switches between modes, as shown in Fig. 15a. On the other hand, the magnitude and direction of the control input vector $\mathbf{u}_{sm}$, used in the modified hybrid closed-loop system (50), change continuously, as depicted in Fig. 16 and Fig. 17, respectively. Therefore, the control input vector $\mathbf{u}_{sm}$ generates smooth robot trajectories, as shown in Fig. 15b.
Figure 16: Magnitude of the control input vector versus time. The red trajectory represents $\|\mathbf{u}\|$, while the blue trajectory represents $\|\mathbf{u}_{sm}\|$.
Figure 17: Plot of the angle between the control input vectors and the positive $x$-axis versus time. The red trajectory represents $\mathbf{atan2v}(\mathbf{u})$, while the blue trajectory represents $\mathbf{atan2v}(\mathbf{u}_{sm})$.
IX Conclusion
We propose a hybrid feedback controller for safe autonomous robot navigation in $n-$dimensional environments with arbitrarily-shaped convex obstacles. These obstacles may have nonsmooth boundaries and large sizes, and can be arbitrarily located, provided they meet certain mild disjointedness requirements, as per Assumption 9. The proposed hybrid controller guarantees global asymptotic stability of the target location in the obstacle-free workspace. The obstacle-avoidance component of the control law relies on the projection of the robot’s center onto the obstacle being avoided, enabling applications in a priori unknown environments, as discussed in Section VII. The proposed hybrid feedback control law generates discontinuous control inputs when switching between modes. A smoothing mechanism has been suggested to overcome this problem in practical applications. Extending our approach to robots, with second-order dynamics, navigating in $n-$dimensional environments with non-convex obstacles would be an interesting future work.
-A Proof of Lemma 1
First, we show that for any $R_{s}>0$, it holds that $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$. As shown in Fig. 18, let us partition $\mathcal{R}_{a}^{i}$ into two mutually exclusive sets as follows:
$\mathcal{R}_{a}^{i}=\mathcal{R}_{1}\cup\mathcal{R}_{2},$ (52)
where the sets $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are given by
$\mathcal{R}_{1}$ =$=\{\mathbf{x}\in\mathcal{R}_{a}^{i}|\eth_{i}^{\mathbf{x}}\cap\left(\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\cap(\mathcal{D}_{r_{a}}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0})))^{\circ}\right)\neq\emptyset\},$
$\mathcal{R}_{2}$ $=\mathcal{R}_{a}^{i}\setminus\mathcal{R}_{1}.$
When $\mathbf{x}\in\mathcal{R}_{1}$, it is straightforward to notice that $\mathbf{x}\in\mathcal{R}_{l}^{i}$, where $\mathcal{R}_{l}^{i}$ is defined in (20). Therefore, we proceed to prove that $\mathbf{x}\in\mathcal{R}_{2}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$. Since $\mathcal{O}_{i}$ is a convex obstacle, when $\mathbf{x}\in\mathcal{R}_{2}$, one has $d(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}),\eth_{i}^{\mathbf{x}})=d(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}),\mathcal{O}_{i})=r_{a}$. Now, if we show that $\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0$ for all $\mathbf{x}\in\mathcal{R}_{2}$, then, as per (20), one can conclude that $\mathbf{x}\in\mathcal{R}_{2}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$.
For $\mathbf{x}\in\mathcal{R}_{2}$, there exists $\mathbf{p}\in\eth_{i}^{\mathbf{x}}$ such that $\mathbf{p}\in\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$. Therefore, it is true that $\mathbf{p}\in\mathcal{H}_{\leq}(\mathbf{x},\mathbf{x})\cap\partial\mathcal{O}_{i}$. Now, let us assume that $\mathbf{x}^{\top}\mathbf{x}_{\pi}<0$, which implies that $\Pi(\mathbf{x},\mathcal{O}_{i})\in\mathcal{H}_{>}(\mathbf{x},\mathbf{x})$. Since $\mathbf{p}\in\partial\mathcal{O}_{i}$ and $\Pi(\mathbf{x},\mathcal{O}_{i})\in\partial\mathcal{O}_{i}$, and obstacle $\mathcal{O}_{i}$ is convex, one has $\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\subset\mathcal{O}_{i}$. Furthermore, since $\mathbf{p}\in\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\cap\mathcal{H}_{\leq}(\mathbf{x},\mathbf{x})$ and $\Pi(\mathbf{x},\mathcal{O}_{i})\in\mathcal{H}_{>}(\mathbf{x},\mathbf{x})$, it is true that $\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\cap\left(\mathcal{B}_{\|\mathbf{x}_{\pi}\|}(\mathbf{x})\right)^{\circ}\neq\emptyset$. This implies that there exists $\mathbf{q}\in\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\cap\mathcal{O}_{i}$ such that $\|\mathbf{x}-\mathbf{q}\|<\|\mathbf{x}_{\pi}\|$, which is a contradiction. Therefore, for $\mathbf{x}\in\mathcal{R}_{2}$, one can conclude that $\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0$, and hence it is proved that for any $R_{s}>0$, $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$.
Next, we prove that if $R_{s}>l_{i}$, then $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$. Notice that the implication $\mathbf{x}\in\mathcal{R}_{a}^{i}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$ has been proved earlier for any $R_{s}>0$. Therefore, we focus on the backward implication and show that when $R_{s}>l_{i}$, $\mathbf{x}\in\mathcal{R}_{l}^{i}\implies\mathbf{x}\in\mathcal{R}_{a}^{i}$. For $\mathbf{x}\in\mathcal{R}_{l}^{i}$, there exists $\mathbf{p}\in\partial\mathcal{O}_{i}$ such that $d(\mathbf{p},\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))\leq r_{a}$, $\mathcal{L}_{s}(\mathbf{x},\mathbf{p})\cap\mathcal{O}_{i}=\mathbf{p}$ and $\mathbf{x}^{\top}(\mathbf{x}-\mathbf{p})\geq 0$. Since $R_{s}>l_{i}$, it is straightforward to notice that $\mathbf{p}\in\eth_{i}^{\mathbf{x}}$. Therefore, according to (19), one can conclude that when $R_{s}>l_{i}$, $\mathbf{x}\in\mathcal{R}_{l}^{i}\implies\mathbf{x}\in\mathcal{R}_{a}^{i}$.
Figure 18: The representation of the sets $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
-B Proof of Lemma 2
Since the target location at the origin $\mathbf{0}$ belongs to the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$, there exists some distance between the target location and the $r_{a}-$dilated obstacle $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$. In other words, since $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ},$ there exists $\bar{\delta}>0$ such that $d(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\bar{\delta}.$ Notice that, according to (24), the set $\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ belongs to the avoidance region associated with obstacle $\mathcal{O}_{i}$ i.e., $\left(\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\right)\subset\mathcal{R}_{a}^{i}$. In addition, as per Lemma 1, $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$, where $\mathcal{R}_{l}^{i}$ is defined in (20). Therefore, it is clear that $\left(\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\right)\subset\mathcal{R}_{l}^{i}$. According to (20), $\Pi(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ does not belong to the set $\mathcal{R}_{l}^{i}$. As a result, one has $d(\mathbf{0},\mathcal{R}_{l}^{i})>\bar{\delta}.$ Hence, it is clear that $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset,$ where $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset.$ Hence, one can set $\bar{\epsilon}\in(0,d(\mathbf{0},\mathcal{R}_{l}^{i})-\bar{\delta}]$ to ensure that $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset$ for all $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for any $\epsilon\in(0,\bar{\epsilon}]$, where the set $\mathcal{ER}^{\mathbf{h}}$ is defined in (28).
-C Proof of Lemma 3
The flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (32) are by construction closed subsets of $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}.$ Hence, condition 1 in Lemma 3 is satisfied.
[…]
-D Proof of Lemma 4
First, we prove that the union of the flow and jump sets covers exactly the obstacle-free state space $\mathcal{K}$. For $m=0$, according to (24) and (25), by construction we have $\mathcal{F}_{0}^{\mathcal{W}}\cup\mathcal{J}_{0}^{\mathcal{W}}=\mathcal{W}_{r_{a}}.$ Similarly, for $m=1$, according to (27) and (30), by construction one has $\mathcal{F}_{1}^{\mathcal{W}}\cup\mathcal{J}_{1}^{\mathcal{W}}=\mathcal{W}_{r_{a}}$. Inspired by [22, Appendix 11], the satisfaction of the following equation:
$\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}},m\in\mathbb{M},$ (53)
along with (26), (31) and (32) implies $\mathcal{F}\cup\mathcal{J}=\mathcal{K}.$
Now, inspired by [22, Appendix 1], for the hybrid closed-loop system (38), with data $\mathcal{H}_{\mathcal{S}}=(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J}),$ define $\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ as the set of all maximal solutions $\xi$ to $\mathcal{H}_{\mathcal{S}}$ with $\xi(0,0)\in\mathcal{K}$. Since $\mathcal{F}\cup\mathcal{J}=\mathcal{K}$, each $\xi\in\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ satisfies rge $\text{rge }\xi\subset\mathcal{K}.$ Additionally, if every maximal solution $\xi\in\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ is complete, then the set $\mathcal{K}$ will be forward invariant [29, Definition 3.13]. Since the hybrid closed-loop system (38) satisfies the hybrid basic conditions, as stated in Lemma 3, one can use [27, Proposition 6.10], to verify the following viability condition:
$\mathbf{F}(\xi)\cap\mathbf{T}_{\mathcal{F}}(\xi)\neq\emptyset,\forall\xi\in\mathcal{F}\setminus\mathcal{J},$ (54)
which will allow us to establish the completeness of the solution $\xi$ to the hybrid closed-loop system (38). In (54), $\mathbf{T}_{\mathcal{F}}(\xi)$ represents the tangent cone to the set $\mathcal{F}$ at $\xi$.
Let $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}\setminus\mathcal{J}$, which implies by virtue of (26), (31) and (32) that $(\mathbf{x},\mathbf{h},\mathbf{a},s)\in(\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}})\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{R}_{\geq 0}$ for some $m\in\mathbb{M}$. For $\mathbf{x}\in(\mathcal{F}_{m}^{\mathcal{W}})^{\circ}\setminus\mathcal{J}_{m}^{\mathcal{W}}$ with $(\mathbf{h},\mathbf{a},m,s)\in\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ the tangent cone $\mathbf{T}_{\mathcal{F}}(\xi)=\mathbb{R}^{n}\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ where the set $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ is given by
$\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})=\begin{cases}\begin{matrix}[l]\mathbb{R}^{n},&\text{if }\mathbf{h}\in(\mathcal{W}_{r_{a}})^{\circ},\\ \mathcal{H}_{\geq}(\mathbf{0},(\mathbf{h}-\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}}))),&\text{if }\mathbf{h}\in\partial\mathcal{W}_{r_{a}},\end{matrix}\end{cases}$ (55)
where for $\mathbf{h}\in\partial\mathcal{W}_{r_{a}}$, the projection $\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}})$ is unique. For $s\in\mathbb{R}_{\geq 0}$, the set $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ is defined as
$\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)=\begin{cases}\begin{matrix}[l]\mathbb{R},&\text{if }s\in\mathbb{R}_{>0},\\ \mathbb{R}_{\geq 0},&\text{if }s=0.\end{matrix}\end{cases}$ (56)
Since, according to (38), $\dot{\mathbf{h}}=\mathbf{0}$ and $\dot{s}=1$, we have $\dot{\mathbf{h}}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\dot{s}\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$, respectively, and (54) holds.
Next, we consider the case where $\xi\in\mathcal{F}\setminus\mathcal{J}$ with $\mathbf{x}\in\partial\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}}$, $m\in\mathbb{M}$. For $m=0$, according to (24) and (25), one has
$\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})\cap\mathcal{R}_{e},$ (57)
and for $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}},$ the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique. Hence, for all $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (58)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively, and $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. Also, according to (12a), for $m=0$, one has $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. According to (23) and (57), for $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$, one can conclude that $\mathbf{u}(\xi)\in\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})$. Moreover, according to (38), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$. Therefore, the viability condition (54) holds for $m=0$.
For $m=1$, according to (27) and (30) one has
$\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}),$ (59)
and for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique and the set $\mathcal{B}_{\lVert\mathbf{x}_{\pi}\rVert}(\mathbf{x})$ intersects with $\partial\mathcal{O}_{\mathcal{W}}$ only at $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ Hence, for all $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (60)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively. Also, according to (12a), for $m=1$, one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a}),\kappa_{r}>0.$ From (16), it follows that $\eta(\mathbf{x})=1$ for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$. As a result, according to (12a) and (13), the control vector is simplified to $\mathbf{u}(\xi)=\kappa_{r}{\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}}.$ Since for any two orthonormal vectors $\hat{\mathbf{h}},\mathbf{a}\in\mathbb{S}^{n-1}$, the matrix $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$ is positive semidefinite, one has $\mathbf{x}_{\pi}^{\intercal}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, for $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, one has $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)^{\intercal}(\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}))\geq 0$. Moreover, according to (38), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$. Hence, the viability condition in (54) holds for $m=1.$
Hence, according to [27, Proposition 6.10], since (54) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a nontrivial solution to $\mathcal{H}$ for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow. They can neither occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded as per definition (30), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$ as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, and would contradict the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ in view of the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times. Furthermore, according to (38), $\mathbf{x}^{+}=\mathbf{x},$ and from the definition of the update law in (34) and (37), it follows immediately that $\mathbf{J}(\mathcal{J})\subset\mathcal{K}$. Hence, the solutions to the hybrid closed-loop system (38) cannot leave $\mathcal{K}$ through jump and, as per [27, Proposition 6.10], all maximal solutions are complete.
-E Proof of Theorem 1
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (38), is immediate from Lemma 4. We next prove the stability of $\mathcal{A}$ using [27, Definition 7.1].
Since $\mathbf{0}\in(\mathcal{W}_{r_{a}})^{\circ}$, there exists $\mu_{1}>0$ such that $\mathcal{B}_{\mu_{1}}(\mathbf{0})\cap(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))^{\circ}=\emptyset.$ According to (24), there exists $\mu_{2}>0$ such that $\mathcal{B}_{\mu_{2}}(\mathbf{0})\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset$. Additionally, as per (27) and (29), there exists $\mu_{3}>0$ such that $\mathcal{B}_{\mu_{3}}(\mathbf{0})\subset\mathcal{J}_{1}^{\mathcal{W}}$. We define the set $\mathcal{S}_{\mu}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{B}_{\mu}(\mathbf{0})\},$ where $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}).$ Notice that for all initial conditions $\xi(0,0)\in\mathcal{S}_{\mu}$, the control input, after at most one jump corresponds to the move-to-target mode and steers $\mathbf{x}$ towards the origin with the control input vector $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. Hence, for each $\mu\in(0,\min\{\mu_{1},\mu_{2},\mu_{3}\}),$ the set $\mathcal{S}_{\mu}$ is forward invariant for the hybrid closed-loop system (38).
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},\bar{s})\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [29, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (38). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
Attractivity: We aim to show that for the proposed hybrid closed-loop system (38), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [29, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
$\underset{(t,j)\in\text{ dom }\xi,t+j\to\infty}{\lim}d(\xi(t,j),\mathcal{A})=\lVert\mathbf{x}(t,j)\rVert=0.$ (61)
The completeness of all maximal solutions to the hybrid closed-loop system (38) follows from Lemma 4. Next, we prove that for all initial condition $\xi(0,0)\in\mathcal{K}$, every complete solution $\xi$ to the hybrid closed-loop system (38), satisfies (61). We consider two cases based on the initial value of the mode indicator variable $m(0,0)$.
Case 1: $m(0,0)=0$. For the hybrid closed-loop system (38), consider a solution $\xi$ initialized in the move-to-target mode. Let us assume $\xi(t_{0},j_{0})\in\mathcal{F}_{0}$ for some $(t_{0},j_{0})\in\text{ dom }\xi,(t_{0},j_{0})\succeq(0,0).$ If $\xi(t,j)\notin\mathcal{J}_{0}$, for all $(t,j)\succeq(t_{0},j_{0})$, then the control input $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=-\kappa_{s}\mathbf{x}$ with $\kappa_{s}>0$ will steer the state $\mathbf{x}$ straight towards the origin. On the other hand, assume that there exists $(t_{1},j_{1})\succeq(t_{0},j_{0})$ such that $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$. Then, according to (34), the control law switches to the obstacle-avoidance mode. As per (24), it is clear that $\mathbf{x}(t_{1},j_{1})\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ for some $i\in\mathbb{I}$. At this instance, according to (34), the proposed navigation algorithm updates the values of the state variables $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{x}(t_{1},j_{1})$, $\mathbf{a}(t_{1},j_{1}+1)\in\mathbf{A}(\mathbf{x}(t_{1},j_{1}))$, $m(t_{1},j_{1}+1)=1$ and $s(t_{1},j_{1}+1)=s(t_{1},j_{1})+1.$ According to (38), $\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$, $\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$ and $m(t_{1},j_{1}+1)=m(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, where $I_{j_{1}+1}=\{t|(t,j_{1}+1)\in\text{ dom }\xi\}$. To proceed with the proof, we need the following lemma:
Lemma 6.
Under Assumption 9, consider a solution $\xi$ to the hybrid closed-loop system (38). If $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$ at some $(t_{1},j_{1})\in\text{ dom }\xi$ such that $\mathbf{x}(t_{1},j_{1})\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for some $i\in\mathbb{I}$, then for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, the following statements hold true:
  1. 1.
    $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$;
  2. 2.
    there exists $t_{2}>t_{1}$ such that $t_{2}<\infty$ and $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1},$
where $\mathbf{h}=\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$ and $\mathbf{a}=\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$.
Proof.
See Appendix -F. ∎
According to Lemma 6, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty$ such that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$. Notice that, according to (27) and (28), one has $\|\mathbf{x}(t_{2},j_{1}+1)\|<\|\mathbf{x}(t_{1},j_{1}+1)\|$. In other words, according to Lemma 6, the proposed navigation algorithm ensures that, at the instance where the control switches from the obstacle-avoidance mode to the move-to-target mode, the origin is closer to the point $\mathbf{x}$ than to the last point where the control switched to the obstacle-avoidance mode. Furthermore, when the control input corresponds to the move-to-target mode, it steers the state $\mathbf{x}$ towards the origin under the influence of control $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Consequently, given that the workspace $\mathcal{W}$ and the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$, are compact, it can be concluded that the solution $\xi(t,j)$ will have a finite number of jumps and will satisfy (61).
Case 2: $m(0,0)=1.$ For the hybrid closed-loop system (38), consider a solution $\xi$ initialized in the obstacle-avoidance mode. Since $m(0,0)=1$, according to (31), $\xi(0,0)\in\mathcal{J}_{1}$. Therefore, according to (37), the control input switches to the move-to-target mode and $m(0,1)=0.$ One can now use arguments similar to the ones used for case 1 to show that the solution $\xi(t,j)$ will have a finite number of jumps and will satisfy (61).
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (38). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (38), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (38) as per [29, Remark 3.5].
-F Proof of Lemma 6
First, we prove that when the control input corresponds to the obstacle-avoidance mode, one has $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. To that end, we make use of Nagumo’s theorem [13, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
$\mathbf{u}(\xi)\in\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x}),$ (62)
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control vector trajectory $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 3, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0$. Since for any two orthonormal vectors $\mathbf{q},\mathbf{s}\in\mathbb{S}^{n-1}$, the matrix $\mathbf{P}(\mathbf{q},\mathbf{s})$ is positive semidefinite, one has $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q},\mathbf{s})\mathbf{p}\geq 0$ for all $\mathbf{p}\in\mathbb{R}^{n}$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (62) holds true.
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})=\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (62) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ (63)
for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Next, we show that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. When the control input corresponds to the obstacle-avoidance mode, it is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$, which, as per (13), can be expressed as a linear combination of the vectors $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$. Note that, according to (34), one has $\mathbf{h}^{\top}\mathbf{a}=0$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as per (14), one has $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Additionally, using (15), one can show that $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Therefore, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, using (63), one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$
for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$ and claim 1 in Lemma 6 is satisfied.
Next, we proceed to prove claim 2 in Lemma 6 which states that when $\xi(t_{1},j_{1}+1)\in\mathcal{F}_{1}$ for some $(t_{1},j_{1}+1)\in\text{ dom }\xi$, the control input steers the state $\xi$ to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$ in finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ with $t_{2}<\infty.$
Let us define the set $\mathcal{O}_{i}^{\mathcal{S}}=\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 19. Since obstacle $\mathcal{O}_{i}$ is convex, the set $\mathcal{O}_{i}^{\mathcal{S}}$ is also convex. As a result, the target location has a unique closest point on the set $\mathcal{O}_{i}^{\mathcal{S}}$, represented by $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}).$ We define a set $\mathcal{LS}$ as follows:
$\mathcal{LS}:=\mathcal{L}(\mathbf{0},\mathbf{0}_{\pi}^{\mathcal{S}})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{H}_{\geq}(\mathbf{0}_{\pi}^{\mathcal{S}},-\mathbf{0}_{\pi}^{\mathcal{S}}),$ (64)
where $\mathbf{0}_{\pi}^{\mathcal{S}}=\Pi(\mathbf{0},\mathcal{O}_{i}^{s})$. Since $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a}),$ the line segment $\mathcal{LS}$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Since $\mathcal{LS}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})=\emptyset$, the line segment $\mathcal{LS}$ also belongs to the exit region $\mathcal{R}_{e}$ (23). Since the hit point $\mathbf{h}$ belongs to $\mathcal{R}_{a}^{i}$, the target location $\mathbf{0}$ is closer to the location $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$ than to the hit point $\mathbf{h}$. Hence, if $\mathbf{0}\notin\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, then for a sufficiently small value of $\bar{\epsilon}$, used in (28), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$ (27). On the other hand, if $\mathbf{0}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, it is straightforward to verify that $\mathcal{LS}\subset\mathcal{S}_{0}$, which, according to (27), implies that $\mathcal{LS}\subset\mathcal{J}_{1}^{\mathcal{W}}$.
Now, if one ensures that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, eventually intersects the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$, then it will imply that $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$, and claim 2 in Lemma 6 will be proven. To that end, let us divide the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 19, into 3 separate subsets as follows:
$\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\mathcal{S}_{3},$ (65)
where the sets $\mathcal{S}_{1},\mathcal{S}_{2}$ and $\mathcal{S}_{3}$ are defined as follows:
$\mathcal{S}_{1}=\mathcal{N}_{\gamma_{a}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$ (66)
$\mathcal{S}_{2}=\left(\mathcal{N}_{\gamma_{s}-\gamma_{a}}(\mathcal{D}_{r_{a}+\gamma_{a}}(\mathcal{O}_{i}))\right)^{\circ}\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$
$\mathcal{S}_{3}=\mathcal{N}_{\gamma-\gamma_{s}}(\mathcal{D}_{r_{a}+\gamma_{s}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$
where $0<\gamma_{a}<\gamma_{s}<\gamma.$
Figure 19: The partition of the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$.
We show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs either to the set $\mathcal{S}_{1}$ or to the set $\mathcal{S}_{3}$, the control eventually steers the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$ Then, we show that for all $\mathbf{x}\in\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi})$. This implies that the state $\mathbf{x}$, which belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ after time $(t_{1},j_{1})$, in the obstacle-avoidance mode around obstacle $\mathcal{O}_{i}$, is always steered to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi})$ and will eventually reach the set $\mathcal{LS}$ at some finite time $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1).$
First, we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ is either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control will eventually steer the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$
When the control input corresponds to the obstacle-avoidance mode and $\mathbf{x}$ belongs to the set $\mathcal{S}_{1}$, the control vector $\mathbf{u}(\xi)$ in (12a) becomes
$\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0.$ (67)
Let $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$ for some $\beta\in[0,\gamma_{a}]$. We know that for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, the tangent cone to the set $\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))$ at $\mathbf{x}$ is given by
$\mathbf{T}_{\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))}(\mathbf{x})=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi}).$
If we show that for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}>0$, then it implies that the control input vector (67) steers $\mathbf{x}$ to the interior of the set $\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 6, ensures that the control input vector (67) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$. To proceed with the proof we require the following fact:
Fact 1: Consider a plane $\mathcal{P}(\mathbf{p},\mathbf{q})$, where $\mathbf{p}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for some $i\in\mathbb{I}$ and $\mathbf{q}\in\mathbf{A}(\mathbf{p})$, where the mapping $\mathbf{A}$ is defined in (35). Then, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$, one has $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\hat{\mathbf{p}},\mathbf{q})\mathbf{x}_{\pi}>0$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$.
Proof.
This proof is by contradiction. First, note that for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\hat{\mathbf{p}},\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ such that $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{p}},\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\hat{\mathbf{p}},\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [30, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))^{\circ}\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.
However, since $\mathbf{q}\in\mathbf{A}(\mathbf{p})$, according to (35), one has $\mathbf{p}_{\pi}\in\mathcal{P}(\mathbf{p},\mathbf{q})$. Therefore, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\subset\mathcal{P}(\mathbf{p},\mathbf{q})$. As a result, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$. This implies that $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})\neq\emptyset$, which is a contradiction. ∎
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, where $\beta\in[0,\gamma_{a}]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}>0$. Therefore, as discussed earlier, the control input vector (67) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
Similarly, when the control input corresponds to the obstacle-avoidance mode and $\mathbf{x}$ belongs to the set $\mathcal{S}_{3}$, the control vector $\mathbf{u}(\xi)$ in (12a) is given by
$\mathbf{u}(\xi)=-\kappa_{r}{\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}},\kappa_{r}>0.$ (68)
Let $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$ for some $\beta\in[\gamma_{s},\gamma]$. We know that for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, the tangent cone to the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$ is given by
$\mathbf{T}_{\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})}(\mathbf{x})=\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x}_{\pi}).$
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, where $\beta\in[\gamma_{s},\gamma]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}>0$. This implies that the control input vector (68) steers $\mathbf{x}$ to the interior of the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 6, ensures that the control input vector (68) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
Finally, we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{H}_{>}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi})$. When the control law corresponds to the obstacle-avoidance mode and the state $\mathbf{x}\in\mathcal{S}_{2}$, according to (12), one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a}),\kappa_{r}>0.$ Note that for all $\mathbf{x}\in\mathcal{S}_{2}$, one has $\eta(\mathbf{x})\in(-1,1)$. Therefore, for every $\mathbf{x}\in\mathcal{S}_{2}$, the vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ can be expressed as a linear combination of the vectors $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}$ given by
$\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})=k_{1}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}+k_{2}\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},$ (69)
where $k_{1}\in\mathbb{R}$ and $k_{2}>0$. Additionally, according to Fact 1, for all $\mathbf{x}\in\mathcal{S}_{2}$, one has $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\neq\mathbf{0}.$ As a result, it can be confirmed that $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})^{\intercal}\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}>0$, when the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2},$ and the proof is complete.
-G Proof of Theorem 2
Forward invariance (Safety): Note that, according to Lemma 5, the hybrid closed-loop system (44) satisfies the hybrid basic conditions. Furthermore, the modified control input vector $\mathbf{u}_{s}(\xi)$, as defined in (42), is obtained by replacing the avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ (13) with $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$ (41) in (12a). Hence, demonstrating that the hybrid closed-loop system (44) satisfies the viability condition, as mentioned in (54), for $m=1$, allows one to employ similar arguments from the proof of Lemma 4 to establish the forward invariance of the set $\mathcal{K}$ for the hybrid closed-loop system (44). In other words, we want to show that for all $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$
$\mathbf{F}_{s}(\xi)\cap\mathbf{T}_{\mathcal{F}_{1}}(\xi)\neq\emptyset.$ (70)
For all $\xi\in\mathcal{K}$ such that $\mathbf{x}\in(\mathcal{F}_{1}^{\mathcal{W}})^{\circ}\setminus\mathcal{J}_{1}$ and $m=1$, the tangent cone $\mathbf{T}_{\mathcal{F}_{1}}(\xi)=\mathbb{R}^{n}\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$, where the sets $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively. Since, according to (44), $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$, the viability condition in (70) holds true for all $\xi\in\mathcal{K}$ such that $\mathbf{x}\in(\mathcal{F}_{1}^{\mathcal{W}})^{\circ}\setminus\mathcal{J}_{1}$ and $m=1$.
Finally, for all $\xi\in\mathcal{K}$ with $\mathbf{x}\in\partial\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ and $m=1,$ the tangent cone $\mathbf{T}_{\mathcal{F}_{1}}(\xi)$ is given by
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (71)
where the sets $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively. According to (42), for $m=1$, $\mathbf{u}_{s}(\xi)=\kappa_{r}\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi},\kappa_{r}>0.$ Note that, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ for each $i\in\mathbb{I}\setminus\{0\}$, one has $\mathbf{x}_{\pi}\in\mathcal{H}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi})$, where $\mathbf{a}\in\mathbb{S}^{n-1}$, and consequently $\mathbf{u}_{s}(\xi)^{\top}\mathbf{x}_{\pi}=0.$ Additionally, according to (44), $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$. Hence, the viability condition in (70) holds true for all $\xi\in\mathcal{K}$ such that $\mathbf{x}\in\partial\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ and $m=1$, and as such it holds true for all $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}.$
Stability: When the mode indicator variable $m=0$, one has $\mathbf{u}_{s}(\xi)=\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0.$ Additionally, the target location at the origin $\mathbf{0}$ belongs to $\mathcal{W}_{r_{a}}^{\circ}$, and the definitions of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$ are the same for (38) and (44). Hence, one can use similar arguments from the proof of Theorem 1 to prove the stability of the target set $\mathcal{A}$.
Attractivity: If we prove that all solutions $\xi$ to the hybrid closed-loop system (44) satisfy Lemma 6, then one can use arguments similar to the ones in the proof of Theorem 1 to prove the attractivity of the target set $\mathcal{A}$ from any point in $\mathcal{K}$ for the hybrid closed-loop system (44). Consequently, we proceed to prove that every solution $\xi$ to the hybrid closed-loop system (44) satisfies Lemma 6.
Since $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$, one has $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ for some $i\in\mathbb{I}\setminus\{0\}$ and $\beta\in[0,\gamma]$, where $\mathbf{h}=\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. First, we show that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$, the tangent cone to the set $\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$ is given by
$\mathbf{T}_{\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})}(\mathbf{x})=\mathcal{H}(\mathbf{0},\mathbf{x}_{\pi}),$ (72)
where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. When the control input vector (42) corresponds to the obstacle-avoidance mode, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$, one has $\mathbf{u}_{s}(\xi)=\kappa_{r}\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, $\kappa_{r}>0$ and $\mathbf{a}=\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. Now, using (41), one can conclude that for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$, $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})^{\top}\mathbf{x}_{\pi}=0$ and $\mathbf{u}_{s}(\xi)\in\mathbf{T}_{\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})}(\mathbf{x})$. Additionally, as per Lemma 5, the control input trajectory $\mathbf{u}(\xi(t,j))$ is continuous when it corresponds to the obstacle-avoidance mode. Therefore, using Nagumo’s theorem [13, Theorem 4.7], one can conclude that for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$
$\mathbf{x}(t,j)\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}),$ (73)
where $\beta\in[0,\gamma]$. Therefore, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Next, we show that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. Since $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$ and obstacle $\mathcal{O}_{i}$ is a sphere, the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ passes through the origin and the center $\mathbf{c}_{i}$ of obstacle $\mathcal{O}_{i}$. As a result, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. In the obstacle-avoidance mode, the control input is given by $\mathbf{u}_{s}(\xi)=\kappa_{r}\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, $\kappa_{r}>0$. According to (41), it is clear that for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})$. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, using (73), one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a}),$ (74)
for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$ and claim 1 in Lemma 6 is satisfied.
We proceed to prove claim 2 in Lemma 6. We know that $\mathbf{h}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ for some $i\in\mathbb{I}\setminus\{0\}$ and $\beta\in[0,\gamma]$. Additionally, according to (73) and (74), it is clear that for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, $\mathbf{x}(t,j)\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, where $\mathbf{a}\in\mathbf{A}_{s}(\mathbf{h})$. Moreover, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, one has $\mathbf{u}_{s}(\xi(t,j))\in\mathcal{H}_{>}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi})$, where $\mathbf{u}_{s}(\xi(t,j))=\kappa_{r}\mathbf{v}_{s}(\mathbf{x}(t,j),\mathbf{h},\mathbf{a}),\kappa_{r}>0$. Therefore, since obstacle $\mathcal{O}_{i}$ is compact, there exists $t_{2}<\infty$ and $t_{2}>t_{1}$ such that $\mathbf{x}(t_{2},j_{1}+1)=\Pi(\mathbf{0},\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))$. Since obstacle $\mathcal{O}_{i}$ is a sphere, it is straightforward to verify that $\Pi(\mathbf{0},\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))\in\mathcal{LS}$, where the set $\mathcal{LS}$ is defined according to (64). If $\mathbf{0}\notin\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, then for a sufficiently small value of $\bar{\epsilon}$, used in (28), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$ (27). On the other hand, if $\mathbf{0}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, it is straightforward to verify that $\mathcal{LS}\subset\mathcal{S}_{0}$, which, according to (27), implies that $\mathcal{LS}\subset\mathcal{J}_{1}^{\mathcal{W}}$. This, according to (27) and (31), implies that there exists $t_{2}\in I_{j_{1}+1}$ such that $t_{2}<\infty$ and $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1},$ and claim 2 in Lemma 6 holds true.
Monotonic decrease of the distance $\|{\mathbf{x}}\|$: Monotonic decrease of $\|{\mathbf{x}}\|$ is trivial in the move-to-target mode, thus, we focus on proving the monotonic decrease in the obstacle-avoidance mode.
Consider a solution $\xi$ to the hybrid closed-loop system (44). Let us assume that there exists $(t_{1},j_{1})\in\text{ dom }\xi$ such that $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$. Therefore, according to (24), (26) and (33), one has $\mathbf{h}(t_{1},j_{1}+1)\in\mathcal{J}_{0}^{\mathcal{W}}\cap\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$, for some $i\in\mathbb{I}\setminus\{0\}$ and $\beta\in[0,\gamma_{s}]$, and $\mathbf{a}(t_{1},j_{1}+1)\in\mathbf{A}_{s}(\mathbf{h}(t_{1},j_{1}+1))$. According to (44), $\mathbf{h}(t,j)=\mathbf{h}(t_{1},j_{1}+1)$ and $\mathbf{a}(t,j)=\mathbf{a}(t_{1},j_{1}+1)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. Let $\mathbf{h}=\mathbf{h}(t_{1},j_{1}+1)=\mathbf{h}(t,j)$ and $\mathbf{a}=\mathbf{a}(t_{1},j_{1}+1)=\mathbf{a}(t,j)$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$. According to Lemma 6, under the control input $\mathbf{u}_{s}(\xi(t,j))$, the state $\mathbf{x}(t,j)$ belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1).$ Moreover, $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$ , where $t_{2}=\underset{t\in I_{j_{1}+1}}{\sup}t.$
If one shows that for all $(t,j)\in([t_{1},t_{2}]\times j_{1}+1),\frac{d}{dt}(\frac{1}{2}\mathbf{x}(t,j)^{\top}\mathbf{x}(t,j))\leq 0$, then it will imply that the control input vector $\mathbf{u}_{s}(\xi(t,j))$ guarantees a monotonic decrease in the distance $\|\mathbf{x}\|$ as the solution $\xi(t,j_{1}+1)$ flows during the interval $I_{j_{1}+1}.$ Notice that for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, one has $\frac{d}{dt}(\frac{1}{2}\mathbf{x}(t,j)^{\top}\mathbf{x}(t,j))=\mathbf{x}(t,j)^{\top}\mathbf{u}_{s}(\xi(t,j))$, where, for $m=1$, $\mathbf{u}_{s}(\xi(t,j))=\kappa_{r}\mathbf{v}_{s}(\mathbf{x}(t,j),\mathbf{h},\mathbf{a})$. Therefore, we proceed to show that $\mathbf{x}(t,j)^{\top}\mathbf{v}_{s}(\mathbf{x}(t,j),\mathbf{h},\mathbf{a})\leq 0$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Let us divide the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ into two mutually exclusive subsets as follows:
$\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})=\mathcal{U}_{1}\cup\mathcal{U}_{2},$ (75)
where
$\mathcal{U}_{1}$ $:=\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i}),$ (76)
$\mathcal{U}_{2}$ $:=\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{<}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{c}_{i}),$
with $\mathbf{c}_{i}$ being the center of obstacle $\mathcal{O}_{i}.$
Figure 20: Geometric representation of the sets $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, defined in (76), for a three-dimensional spherical obstacle $\mathcal{O}_{i}$.
Since $\mathbf{a}$ and $\mathbf{h}$ are chosen as per (33), and obstacle $\mathcal{O}_{i}$ is a sphere, the plane $\mathcal{P}(\mathbf{h},\mathbf{a})$ intersects both the center $\mathbf{c}_{i}$ of obstacle $\mathcal{O}_{i}$ and the target location at the origin, as shown in Fig. 20. As a result, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, one has $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ and $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})$. Moreover, since obstacle $\mathcal{O}_{i}$ is a sphere, for all $\mathbf{x}\in\mathcal{U}_{1}$, it is true that $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x})$. Therefore, for all $\mathbf{x}\in\mathcal{U}_{1}$, one has $\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x})$, as illustrated in Fig. 21. As a result, for all $\mathbf{x}\in\mathcal{U}_{1}$, one can conclude that $\mathbf{x}^{\top}\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})\leq 0$. Now, if one shows that $\mathbf{x}(t,j)\in\mathcal{U}_{1}$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, then it implies that $\mathbf{x}(t,j)^{\top}\mathbf{u}_{s}(\xi(t,j))\leq 0$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$
Figure 21: Workspace scenario illustrating $\mathbf{x}_{\pi}\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{R}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x})\implies\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})\in\mathcal{P}(\mathbf{h},\mathbf{a})\cap\mathcal{H}_{\leq}(\mathbf{0},\mathbf{x})$.
Note that, when $\xi(t_{1},j_{1})\in\mathcal{J}_{0}$ with $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, the choice of the unit vector $\mathbf{a}$, as per (35), ensures that $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{U}_{1}.$ We know that, under the control input $\mathbf{u}_{s}(\xi(t,j))$, the state $\mathbf{x}(t,j)$ belongs to the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{h},\mathbf{a})$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, where $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}),$ for some $\beta\in[0,\gamma_{s}]$. Since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{U}_{1}$ and $\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})^{\top}\mathbf{x}\leq 0,\forall\mathbf{x}\in\mathcal{U}_{1}$, under the control input $\mathbf{u}_{s}(\xi)=\kappa_{r}\mathbf{v}_{s}(\mathbf{x},\mathbf{h},\mathbf{a})$, $\mathbf{x}$ can enter in the set $\mathcal{U}_{2}$ only from the set $\mathcal{LS}$, where $\mathcal{LS}$ is defined in (64) and is depicted in Fig. 20. Therefore, there exists $(t_{2},j_{1}+1)\succ(t_{1},j_{1}+1)$ such that $\mathbf{x}(t_{2},j_{1}+1)=\mathcal{LS}\subset\mathcal{U}_{1}$. Moreover, as proved earlier, for a sufficiently small value of $\bar{\epsilon}$, used in (28), one can guarantee that $\mathcal{LS}\subset\mathcal{J}_{1}^{\mathcal{W}}$, where the set $\mathcal{J}_{1}^{\mathcal{W}}$ is defined in (27). Hence, at $(t_{2},j_{1}+1)$, one has $\xi(t_{2},j_{1}+1)\in\mathcal{J}_{1}$, which implies that $t_{2}=\underset{t\in I_{j_{1}+1}}{\sup}t.$ As a result, for all time $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, one has $\mathbf{x}(t,j)\in\mathcal{U}_{1}$, and the proof is complete.